### v1

# CayleyPy RL: Pathfinding and Reinforcement Learning on Cayley Graphs

Abstract.
This paper is the second in a series of studies on developing efficient artificial intelligence-based approaches to pathfinding on extremely large graphs (e.g. $10^{70}$ nodes) with a focus on Cayley graphs and mathematical applications. The open-source CayleyPy project is a central component of our research. The present paper proposes a novel combination of a reinforcement learning approach with a more direct diffusion distance approach from the first paper. Our analysis includes benchmarking various choices for the key building blocks of the approach: architectures of the neural network, generators for the random walks and beam search pathfinding. We compared these methods against the classical computer algebra system GAP, demonstrating that they ”overcome the GAP” for the considered examples. As a particular mathematical application we examine the Cayley graph of the symmetric group with cyclic shift and transposition generators. We provide strong support for the OEIS-A186783 conjecture that the diameter is equal to n(n-1)/2 by machine learning and mathematical methods. We identify the conjectured longest element and generate its decomposition of the desired length. We prove a diameter lower bound of n(n-1)/2-n/2 and an upper bound of n(n-1)/2+ 3n by presenting the algorithm with given complexity. We also present several conjectures motivated by numerical experiments, including observations on the central limit phenomenon (with growth approximated by a Gumbel distribution), the uniform distribution for the spectrum of the graph, and a numerical study of sorting networks. To stimulate crowdsourcing activity, we create challenges on the Kaggle platform and invite contributions to improve and benchmark approaches on Cayley graph pathfinding and other tasks.
Key words and phrases:
Machine learning, reinforcement learning, Cayley graphs
1. Introduction
1.1. Broader context
Deep learning has revolutionized various fields of research and was recognized by several Nobel Prizes in 2024. In recent years, machine learning has been emerging as ”a tool in theoretical science” [Douglas22], leading to several noteworthy applications to mathematical problems: [Charton19, Davies21, Bao23, Romera24, Coates23, Hayat25, Williamson24, Gukov24, Kohli25].
The paper presents advancements in applying artificial intelligence methods to the mathematical problem of Cayley graph pathfinding, which is equivalent to decomposing a group element into a product of generators. It is the second work in a series devoted to the CayleyPy project to develop efficient machine learning-based approaches and an open source library to deal with extra large graphs (e.g. $10^{70}$) with a focus on finite Cayley graphs and mathematical applications.
From a broader perspective, pathfinding is a specific case of the planning problem, where one must determine a sequence of actions to transition between two given states. Similar challenges arise in robotics and games like chess or Go, where the objective is to plan moves that lead from the initial position to a winning position. Mathematically, such problems are modeled as pathfinding on graphs (state transition graphs), where nodes represent possible states, and edges correspond to state transitions based on actions or moves. The planning task then reduces to finding a path from a given initial node to one or more target nodes. The breakthrough works AlphaGo and AlphaZero [AlphaGo16], [AlphaZero17] by Google DeepMind have demonstrated that machine learning can significantly outperform previously known methods. They serve as both a prototype and an inspiration for many developments, including ours. In a nutshell, the method consists of two main components: a neural network trained to suggest moves from any given state, and a graph pathfinding algorithm that helps to find a path even when the neural network’s predictions are not sufficiently precise. The core idea is similar across most approaches, but achieving optimal performance for each task requires analyzing and benchmarking various neural networks and graph pathfinding methods. One of the goals of this work is to present such an analysis for some particular Cayley graphs.
1.2. Present paper: approach, benchmarks, mathematical theorems and conjectures
Figure 1. LRX Cayley graph for $S_{5}$
The paper presents both AI contributions and mathematical results and conjectures. Here, we work with a particular Cayley graph defined by the L - left cyclic shift, R - right cyclic shift and X - eXchange of positions $0,1$, it will be denoted LRX following OEIS-A186783. Its image for $S_{5}$ is shown on figure 1.
Developments of AI methodology. We propose a modified deep Q-learning (DQN) method for the case of the graph pathfinding task which resolves the so-called sparse reward problem of the standard DQN approach for graph pathfinding. The proposal combines DQN with the more simple diffusion distance approach from our previous paper  [Chervov25]. An analysis and benchmarks of these approaches are presented, including tests of various neural network architectures such as transformers, convolutional networks, and perceptrons. It is demonstrated that these methods allow finding paths for $S_{n}$ up to $n$ around 30 - for the particular class of generators (LRX) which are in the main focus of the present paper. We demonstrate that they ”overcome the GAP” - outperform classical computer algebra system GAP on the same task which can work only up to $n$ around 20.
Efficient pytorch implementation. We pay significant attention to develop an original highly optimized code, the current version is specific to Cayley graphs of permutation groups (matrix groups will be supported later). We propose many technical solutions for fast parallel processing of large amounts of permutations, including product computations, efficient hashing, and extracting unique states. The code works on both CPU and GPU without modification.
One line of code bringing orders of magnitudes improvement. We report a curious finding - a single-line modification in the code of our main graph pathfinding module (beam search), dramatically improves performance. This enhancement increases the feasible pathfinding size from symmetric groups $S_{n}$ with $n$ to $n$ around 100 (and apparently even further with more hardware). Unfortunately, this improvement is currently specific to the particular LRX Cayley graphs considered in this paper, but we hope it can be generalized.
Mathematical contributions.
LRX Cayley graphs. Diameter conjecture OEIS-A186783. In the present paper, we focus on a specific family of Cayley graphs for the symmetric groups $S_{n}$ those generated by the two cyclic shifts (L) - left, (R) - right and the transposition X=$(0,1)$ of the first two positions. Surprisingly, according to OEIS-A186783, it is an open conjecture that the diameter of the Cayley graph is $n(n-1)/2$ for $S_{n}$. These generators and Cayley graphs will be denoted LRX as in OEIS.
The longest element. By implementing an efficient brute-force breadth-first search algorithm, we traverse the full LRX graphs for $n\leq 14$, confirming that the longest element has a length of $n(n-1)/2$, as predicted. Moreover, we observe that this element is always unique and follows a clear pattern - it is a product of transpositions: $(0,1)(n-1,2)(n-2,3)(n-3,4)...$, so we expect this for all $n$.
AI and crowd-sourcing to support the conjecture. For that particular element, we have launched various versions of our pipelines and have consistently found that the decomposition is never shorter than $n(n-1)/2$. We organized a challenge on the Kaggle platform giving that particular element for $n=2...200$ and requesting decompositions of the shortest possible length. Once again, no participant found a decomposition shorter than $n(n-1)/2$, thereby confirming the expectation.
Rigorous lower and upper bounds. We prove lower bound $n(n-1)/2-n/2-1$ by a combinatorial argument (improving $n(n-1)/3$ [Babai89]); afterwards, we prove the diameter upper bound $n(n-1)/2+3n$ by the algorithm with such complexity (improving $3/2n^{2}$  [Kuppili20]). We also develop another algorithm which empirically shows better complexity: $n(n-1)/2+n/2$, but lacks the rigorous proof for such an estimate. Finally, we present explicit decomposition for the conjecturally longest element of the desired length $n(n-1)/2$.
Central limit phenomena for growth - Gumbel distribution. Computations suggest the following conjecture: for large $n$, the growth function of LRX graphs follows an asymmetric Gumbel distribution. This is in the vein of field-shaping works by P. Diaconis et. al. [Diaconis77, Diaconis88, Diaconis17], who demonstrated that the growth of the neighbor transposition graph and many related statistics asymptotically follow the Gaussian normal distribution.
We study numerically the spectrum of LRX graph and observe surprisingly uniform distribution of eigenvalues, we also analyze statistics of the paths from the longest element to the identity (”sorting networks”) we observe numerically that pattern differs from the famous ”sine curves” for Cayley graph of neighbor transpositions [Virág]; numerical analysis of random walks shows consistency with results of Diaconis [Diaconis93] on mixing time for LRX Cayley graph.
We also present similar analysis for the Schreier coset graph for the action of the LRX generators on binary strings consisting of $n/2$ zeros and ones (for even $n$).
In summary, our analysis demonstrates that computational experiments, especially those involving AI, can be very helpful in advancing mathematical research on Cayley graphs. AI methods significantly outperform classical methods of computer algebra systems such as GAP, performing well for $S_{n}$ with $n$ around 30, and with a small addition of prior knowledge up to 90, while classical methods perform well only up to 20.
1.3. Finite Cayley graphs: pathfinding techniques, results and open problems
Let us briefly summarize existing techniques for Cayley graphs pathfinding, and remind more general context on Cayley graphs research.
Optimal pathfinding is typically NP-hard. Finding the shortest paths on generic finite Cayley graphs is an NP-hard problem [Goldreich81] (even P-space complete [Jerrum85]). As is the case for many specific group families, such as NxNxN Rubik’s Cube groups [Demaine17] and others [Bulteau15]. For particular groups it is hard to develop practical optimal algorithm e.g. standard 3x3x3 Rubik’s cube the first optimal algorithm has been proposed only 1997 [Korf97]. It required 192 Gigabytes of precomputed data and 12-24 hours to solve single cube (methods were later improved  [^1] M.Reid: https://www.cflmath.com/Rubik/optimal_solver.html, H.Kociemba 2021: https://pypi.org/project/RubikOptimal/, Andrew Skalski: https://github.com/Voltara/vcubem , [Rokicki14] to several cubes per second). To develop practical optimal solver for 4x4x4 Rubik’s cube - is a challenge.
Non-optimal pathfinding: Schreier-Sims algorithm, GAP implementation, non-effectivity for large groups. The Schreier-Sims algorithm [Sims70],[Knuth91] (or its randomized versions) is the method typically used for decomposing group element into product of generators for permutation groups. It is implemented in GAP computer algebra system. However its outputs ”are usually exponentially long” [Shamir89] and that forbids practical computations for large groups (typically the sizes beyond $10^{30}-10^{40}$ are out of reach - examples and analysis is given below).
Non-optimal pathfinding: ”small support” Kaggle Santa 2023 methods (generality issue). The methods which can deal with extremely large groups (e.g. size $>10^{1000}$ like 33x33x33 Rubik’s cube) were demonstrated by top participants of the recent challenge organized on the Kaggle platform - Santa 2023. However the generality of that approach is unclear.That is contrast to AI methodology which is from the beginning is general not even restricted to Cayley graphs. The idea of ”Santa methods” is quite noteworthy: one should first find ”small support elements” achievable from original generators and take them as new generators, then one simply runs beam search with Hamming distance heuristics to find path. Method becomes effective since for small support generators Hamming distance and true graph distance are quite related by obvious reason. ”Small support” are those permutations which change only small number of positions. They are well-known for Rubik’s cube solvers - ”commutator” or ”Pif-Paf” moves [Mulholland16], M.Brandenburg post. In mathematics small support elements play an important role in deep works by Babai, Seress, Helfgott, et.al. [Babai88, Babai04, Seress14, Helfgott14, Helfgott19] (slides: A.Seress,H.Helfgott ), however it is unclear whether one can effectively find small support elements from given generators, thus making unclear the generality of ”small support methods”.
Non-optimal pathfinding: algorithms for specific families of generators. For some particular choices of generators there are algorithmic solvers which can effectively find non-optimal solution. It seems unlikely that one can produce similar algorithms for any generators. The prototypical example is bubble sort algorithm, which solves the problem for the neighbor transpositions generators of $S_{n}$. Algorithmic solvers exist for Rubik’s cube of arbitrary sizes and for other puzzles related groups. The other families of generators and algorithms are important in bioinformatics for estimation of the evolutionary distance. These generators are related to flips of subsequences - see e.g. [Pevzner95, Pevzner99, Bulteau19]. Related example is pancake sorting (or prefix sorting) problem with the first algorithm proposed by Bill Gates et.al. [Gates79].
AI-methods for pathfinding. To the best of our knowledge there is no systematic effort for Cayley graph pathfinding with AI methods except our CayleyPy project. However for some particular cases for example Rubik’s cube there exists a number of approaches, e.g. based on genetic algorithms: [Świta23].The most notable are: DeepCube [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] - the first AI based successful solver for 3x3x3 cube (website), the second one: EfficientCube [Takano21], some other: [TakanBrunetto17, Johnson21, Amrutha22, Noever21, Chasmai21, Bedaywi23, Pan21] which not achieve complete solution. Noteworthy idea [Pan21] is to combine neural networks with representations theory of the symmetric group — neural net predicts the coefficients of the non-abelian Fourier transform for the distance function. The rationale is: observed sparsity (bandlimitedness) of Fourier transform of the common distance functions on $S_{n}$  [Swan17]. The recent seminal achievement from S.Gukov’s team:  [Gukov24] creates effective AI based pathfiding approach for infinite group of Andrews-Curtis moves and resolves the Akbulut-Kirby conjecture remained opened for 39 years. AI-methods are also used for pathfinding in the context of planning movements in obstacle-rich environments and road networks it is related field, but with different focus [Leskovec22, Yakovlev23]
Cayley graphs: applications, diameters, random walks, open conjectures. Let us give a more general context on Cayley graph research. Cayley graphs are fundamental in group theory [Gromov93],[Tao15], and have various applications: bioinformatics  [Pevzner95, Pevzner99, Bulteau19]; processors interconnection networks [Akers89, Cooperman91, Heydemann97];in coding theory coding theory and cryptography [Wigderson06, Zémor94, Petit11]; in quantum computing [Kohli24, Sarkar24, Vidick23, Acevedo06, Gromada22], etc.
Two focuses of pure mathematical research on finite Cayley graphs are related to diameters estimation and the behavior of random walks. In the case of the symmetric group $S_{n}$ there are two open conjectures which are easy to formulate, but somewhat representative for the field.
Babai-like conjecture: diameter of $S_{n}$ is not more than $O(n^{2})$ for any choices of generators (see e.g. H.Helfgott’s surveys:  [Helfgott14],[Helfgott19],[Helfgott15] ;
Diaconis  [Diaconis13] conjecture: the mixing time for random walks is not more than $O(n^{3}log(n))$ (again for any choices of generators).
The first one can be thought as a particular case of the Babai conjecture [Babai92], [Tao15] which predicts that diameter of simple groups is not as large as one may expect: $O(log^{c}|G|)$ for some absolute $c$ — in contrast to abelian groups where it can be $O(|G|)$. Thus conjecture represents an intriguing connection between algebraic properties of a group (which does not depend on the choice of generators) - being simple and the property of Cayley graphs - having short diameter (which is generator dependent property).
Estimating diameters of Cayley graphs is a hard problem. One of its applications is latency estimation for processor interconnection problem [Akers89],[Cooperman91], [Heydemann97] - since diameter corresponds to maximal latency. It is also of interest for community of puzzles enthusiasts - like Rubik’s cube. Diameter is also called a ”God’s number” - the number of moves the worst configuration can be solved by the best (optimal) algorithm. It is widely unknown for most of the puzzles, its determination for the Rubik’s cube 3x3x3 for the standard choices of generators required significant efforts and computational resources [Rokicki14]. It is not known precisely for higher cubes, neither for 3x3x3 Rubik’s with STM, ATM metrics.A precise conjectures on 4x4x4 and 5x5x5 Rubik’s cube diameters appeared recently [Hirata24], collections of conjectures and estimates for many puzzles can be found on speedsolving website.
As we will discuss below novel AI methods might provide insights via computational experiments to such kind of questions.
2. Reinforcement learning (RL) and graphs pathfiding
2.1. Reminders: RL, sparse reward problem, Bellman equation for graph distances, Q-algorithm, pathfinding
RL in brief. Reinforcement learning is an important direction in AI that differs from supervised and unsupervised learning in the following ways: it focuses on optimizing cumulative rewards, and it requires the agent to, in a sense, ”create its own training set”. The agent must choose the best strategies to ”explore the space” and select training examples on its own, in contrast to supervised learning, where the training set is provided in advance.
RL ”equivalence” to graph pathfinding. The setup of reinforcement learning is very similar to the setup of graph pathfinding: states - are nodes of the graph, actions corresponds to edges/transitions to neighbor states by these edges, rewards corresponds to weights on the edges. The cumulative reward corresponds to length of paths (with respect to edges weights). Typically in graph theory shortest paths are of interest - that means ”reward” is minimized, not maximized as in RL, so it is better to use term not the ”reward”, but ”penalty”.
Unweighted graphs and sparse reward problem. The sparse reward problem is a well-known challenge in reinforcement learning (RL), and it specifically arises in pathfinding on Cayley graphs and, more generally, on any unweighted graph, which makes standard RL approaches perform poorly. The sparse reward problem occurs when rewards are given very rarely, making it difficult to generate data for training. This situation is particularly prevalent in graph pathfinding on large graphs, where the task is to find a path to a single node out of trillions of others. A reward is only given once the desired node is reached, but it is practically impossible to find it with a ”cold start” through random exploration. It’s important to note that the condition of all weights being equal to one plays a significant role. If the weights were different, one could reasonably adopt a strategy of selecting the step with the best local reward. However, when all weights are identical, there is no local preference to guide the agent in making the right move. There are several strategies to address this issue, and below, we propose a new approach.
Bellman value equation is equation for distance functions. The value function (or ”position estimation”) in RL exactly corresponds to the distance functions on graphs. To make that analogy more precise: consider an unweighted graph $G$ (all edges weights are equal to 1), choose some node $e$ (e.g. identity of a group in Cayley graph), consider a function $d(g)$ the distance of the shortest path from $g$ to $e$ (it is ”word metric” in group theory). Then there is a simple relation:
$d(g)$ $=1+\min_{n:~{}neigbours~{}of~{}g}d(n)\text{~{}~{}~{}~{}~{}~{}~{}~{}~{} Bellman equation}$ (2.1)
$d(e)$ $=0\text{~{}~{}~{}boundary condition}$ (2.2)
This relation is simple, but it is fundamental - it is a particular case of the Bellman’s equation. The properties can be summarized in the following simple proposition, which is easy to see inductively considering neghbor nodes to $e$, then neighbors of neighbors and so on.
Proposition 2.1.
For any finite connected graph $G$ the only solution to the pair of equations above are given by the distance function (length of the shortest path) to the node $e$. Moreover the only solutions to the Bellman’s equation (without boundary condition) correspond to distance to some sets of nodes in $G$ (up to adding a constant). That set can be recovered as minimums of $d(g)$.
Sense of Bellman equation - local condition for globally (cumulatively) defined object. The Bellman equation is equation for the length of the shortest path - that is globally/cumulatively defined object, but the condition itself is local - only at the neighbor of each point.
Idea of Q-learning and deep Q-learning algorithms for value/distance function. These algorithms provide a way to compute solutions of the Bellman equation, which means to find the distance functions on graphs. One starts with some initialization for $d(g)$, then compute next iteration:
$d_{k}(g)=1+\min_{n:~{}neigbours~{}of~{}g}d_{k-1}(n),d_{k}(e)=0$ by applying the right hand side of the equations 2.1. There are theoretical guarantees that such a process will converge to the solution of the Bellman equation (and boundary condition).
Idea of deep Q-learning algorithm. The deep Q-learning is used in the situations where the graph is large (can be of google size). One assumes that for each node of graph there is a feature vector attached. One considers a neural network which takes a feature vector as an input and should output the approximate solution for Bellman equations. The training process is organized as follows. Each epoch one selects some subset of nodes and from current neural network $d(g)$ computes the right hand side of the equations 2.1: $t(g)=1+\min_{n:~{}neigbours~{}of~{}g}d(n),t(e)=0$, denoting $t(g)$ as a result of that computation. Afterwards one runs gradient descent to minimize the error between $d(g)$ and $t(g)$. If the error would be zero that means that $d(g)$ solves the Bellman equations. And thus in graph theory setup it produces the distance function (”word metric” in the case of Cayley graphs). Typically the choice of the training subset of nodes is made with the help of walking over the graph combining both random steps and steps guided by the model. We will discuss that in details below.
Graph path finding. To find a path from a given node $g$ to the selected node $e$, one needs to run a greedy search algorithm. Starting from node $g$, the algorithm inspects all its neighbors, evaluates the distance function for each of them, and then chooses the node with the minimum value of the distance function, moving to that node. The process is repeated until the destination node is found or the step limit is exceeded. If the distance function accurately reflects the true distance on the graph (i.e., the length of the shortest path), this algorithm will find the shortest path in an optimal way, performing the least number of operations. However, in the case of deep Q-learning, the predictions of the neural network are typically not very precise, and the greedy algorithm can get stuck in local minima. To overcome this challenge, more advanced graph search algorithms, such as beam search or A-star, are used. This will be discussed further below.
2.2. Diffusion distance pathfinding - CayleyPy-1
Let us briefly remind the novel pathfinding approach proposed in our previous paper [Chervov25]. Its efficiency has been demonstrated on the Rubik’s cube where it significantly outperforms all possible competitors providing shorter solutions than any existing approach for 3x3x3,4x4x4,5x5x5 cubes. It is also more computationally efficient and in some sense it is more simple.
The idea of the approach is to work not with the true distance on the graph, but with the diffusion distance which might be thought as a length of the random path. The diffusion distance is easy to estimate by the number of steps of random walks. Generation of random walks is much more fast than computations of the neural networks predictions (the only way to generate target in Q-learning), thus generation of the training data is much more efficient than in Q-learning approach, though the training data might be of weaker quality.
The assumption is that for each node of a graphs there is feature vector attached. For example for permutation groups it is a vector representing a permutation $(p_{0},p_{1},p_{2},...p_{n-1}),p:i\to p_{i}$ We also assume that there is a selected node (e.g. identity of the group) where we need to find a path from any other node.
Core of the approach consists of the three steps:
  1. (1)
    Generation of training set by random walks. Generate $N$ random walks trajectories all starting from a selected node $e$ (identity of a group). Each random walk trajectory consists of up to $K_{\text{max}}$ steps, where $N$ and $K_{\text{max}}$ are integer parameters of the method. For some nodes encountered during the random walks, we store a set of pairs $(v,k)$, where $v$ represents the vector corresponding to the node and $k$ is the number of steps required to reach it via the random walk.
  2. (2)
    Generation of training set by random walks. The generated set of pairs $(v,k)$ serves as the training set for the neural network. Specifically, $v$ serves as the ’feature vector’ (the input for the neural network), and $k$ represents the ’target’ (the output that the network needs to predict). Thus, the neural network’s predictions for a given node $v$ estimate the diffusion distance from $v$ to the selected destination node.
  3. (3)
    Graph search guided by neural network. Beam search. This step finds a path from a given node to the destination node. The neural network provides heuristics on where to make the next steps, while the graph pathfinding technique compensates for any possible incorrectness in the neural network predictions. The beam search pathfinding method is quite simple, but has proven to be the most effective for us and works as follows. Fix a positive integer $W$ a parameter known as the ”beam width” (or ”beam size”). Starting from a given node, we take all its neighboring nodes and compute the neural network predictions for all of them. We then select the $W$ nodes that are closest to the destination according to the neural network (i.e., the predictions have smaller values). For these selected $W$ nodes, we take their neighbors, drop duplicates, and again compute the neural network predictions, choosing the top $W$ nodes with the best (i.e., minimal) predictions. This process is repeated until the destination node is found (or limit of steps is exceeded).
The last step is common for all methods - based on Q-learning or diffusion distance. One uses a neural network predictions as heuristic function and uses graph search algorithms to find a path. We are strongly in favor of beam search, while DeepCube team [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] uses A-star algorithm and develops its modifications.
2.3. Beam search effectivity - toy model of research community
To some extent, the success of the proposed scheme is due to beam search and our efficient implementation supporting large beams. We started with a modified greedy search, then moved to Metropolis and A-star, before finally finding relief in beam search. The larger the beam size, the more effective beam search becomes (probability to find path increases and paths become shorter). To maximize beam sizes, we developed an original and efficient PyTorch implementation capable of supporting beam sizes with millions of nodes and beyond (basic, advanced versions). On one hand, beam search is arguably one of the simplest methods. On the other hand, there are several key aspects worth highlighting.
The primary reason we need beam search is to avoid the problem of local minima. Greedy search can get stuck in them, whereas beam search can bypass them entirely if the beam size is larger than the depth of the local minima. It’s similar to how an elephant wouldn’t even notice small dips in the terrain, while for an ant, they could be insurmountable obstacles.
Another important aspect of beam search is that increasing the beam size not only linearly increases computation but also linearly increases memory consumption. On one hand, typically memory - not computational time - is the main bottleneck of the method. On the other hand, having the largest possible beam size is crucial for its performance. Moreover, there is no simple way to reproduce the results of beam search with a beam size of $B+1$ using any algorithm that only consumes memory equivalent to beam size $B$.
To explain this point, we use an analogy with the work of a research community.
Imagine two researchers working together on a creative problem (a beam size of 2). The key factor is the exchange of ideas—if one researcher makes a breakthrough, they share it with the other, allowing both to continue their search from the breakthrough point. Without this exchange, the second researcher might get stuck in a dead end. This sharing of ideas helps avoid wasting time on unproductive directions. The mechanism described above explains why increasing the beam width is so effective—it acts as an exchange of ideas, focusing on breakthroughs rather than following unproductive paths. Essentially, this is what beam search does: it selects the top nodes across the entire beam neighborhood rather than picking the best few for each individual node. This means that if the neighbors of one node are significantly closer to the destination (analogous to a breakthrough), those nodes are retained in the beam, while the neighbors of less successful nodes are discarded. In other words, the descendants of some nodes do not survive into the next generation. A kind of natural selection process.
From the above, it is clear that the size of a research community matters. Reducing the community size by half will not simply cut the research output in half — it will decrease it exponentially. That aspect is different from the more routine work - cutting it twice output will be cut twice. The same principle applies to beam search: the beam size is crucial, and there is no simple way to achieve the performance of a beam size $B$ using a smaller beam.
This logic also explains why evolutionary or nature-inspired global optimization methods can sometimes be successful. They rely on the same fundamental mechanism as beam search but offer better control over the diversity within the beam. In some cases, it is important not to immediately jump to breakthroughs made by others but to continue exploring independent directions, as this might lead to even greater discoveries. For example, if medical doctors would stop doing their job to start follow breakthroughs in mathematics, it would lead to disastrous results — diversity is crucial in the long run. The balance between ”everyone follows the latest breakthrough” (as in beam search) and maintaining diversity is delicate. There is no universal solution, as different tasks require different trade-offs. Exploring various methods on Cayley graph pathfinding tasks is an interesting direction.
2.4. Novel method - combining diffusion distance with deep Q-learning
Rationale. Why we need that? The diffusion distance approach described above is quite simple and efficient however there is practical and theoretical disadvantage. The practical side consists of somewhat surprising phenomena that enlarging the training set (generating more random walks) does not always improve the quality of the model - such stagnation was observed for Rubik’s cube cases in our previous paper and it is even more prominent for LRX graphs. It is actually related to the quite clear theoretical fact: diffusion distance is not monotonically dependent on the true distance - one node can be further than another for the true distance, while oppositely ranged for the diffusion distance. (That problem is not only for random walks estimation, but for the diffusion distance itself (i.e. limit of infinite data - computed theoretically (analysis) ). So diffusion distance is easy to estimate, but it is not precise. On the other hand DQN and Bellman equation allows to approximate the true distance, but for the price of more heavy computations. The idea of the proposal below is to combine less precise but light weight diffusion distance with more heavy and more precise Bellman equation approach. And to avoid the sparse reward problem.
Our proposal looks as follows.
Part 1. Warm-up diffusion distance training. Use the diffusion distance approach for preliminary training the neural network. I.e. apply the steps 1,2 above: generate random walks to create training data, and train neural network to predict number of steps of random work.
Part 2. Modified DQN training. The second phase will relies on the Bellman equation strategy with mild modifications, but on somewhat different space exploration strategy comparing to conventional DQN - borrowed from the diffusion distance approach:
Step 1. Generate $N$ non-backtracking random walks trajectories all starting from a selected node $e$ (identity of a group). Each random walk trajectory consists of up to $K_{\text{max}}$ steps, where $N$ and $K_{\text{max}}$ are integer parameters of the method.
Step 2. For the obtained nodes compute the Bellman equation predictions $t(g)=1+\min_{n:~{}neigbours~{}of~{}g}d(n),t(e)=0$ in the standard way.
Step 3. Clip the predictions by the number of steps of the random work since true distance is always smaller than number of steps. (Clip also predictions negative values).
Step 4. Run gradient descent for the neural network to minimize the loss function between clipped $t(g)$ (target) and neural network predictions.
Part 3. Graph search algorithm - beam search with neural network as an heuristics. After the neural network is trained we utilize the same approach to find a path by beam search and neural network as in our previous paper.
The advantages of the proposed method are the following: there is no sparse reward problem - since we always start from the selected state, it always allows us to clip overestimated predictions, the warm-up phase allows to avoid long initialization process until when targets obtained by Bellman equation random since network starts from random weights - in our approach targets are always meaningful - first they are steps of random walks, at the second phase neural by itself produces meaningful targets since it was preliminary trained. We also suggest to use non backtracking random walks by rather clear reason - number of steps for non-backtracking is more similar to true distance. Conceptually that is related to the well-known phenomena - ”Non-backtracking random walks mix faster” [Alon07].
Benchmarks and analysis. Below we present the analysis of that proposal. We observe stable improvement of the method above comparing to diffusion distance approach, unfortunately the effect is not so big.
The first experiment is concerned not large groups, where one can compute by brute force the true graph distance and compare it with the predictions of the neural network. We see that for small sizes of the group methods learn practically ideally the distance on a graph, however for large groups correlation becomes not so perfect. Additional training by DQN always improves the preliminary warm-up training by diffusion distance procedure.
$n$ epochs: 5+0 epochs: 50+0 epochs: 5+50 epochs: 50+50
6 0.955 0.99 0.99 0.99
8 0.954 0.969 0.98 0.982
10 0.883 0.923 0.939 0.938
Table 1. n - permutation length. Spearman correlation of the true distance (obtained by brute force on million states) vs predictions of the neural networks trained by different methods. Conclusion Additional training with DQN improves results. Epochs A+B denotes: A - training epochs in warm-up diffusion distance part, B - epochs in DQN part.
The second experiments concerns larger group $S_{28}$ with LRX generators. We train the model and search for path by the beam search. The length of the path is the measure of performance. Experiment confirms that additional training by DQN improves the results. The combinatation of the 30+200 epochs training with 30-warmup and 200 - second stage (DQN) - is one the best which we observed for that case. Average length is 402, best is 382 which is not far from the expected ideal length 378. (In all experiments we searched path from the same element - conjecturally the longest element for LRX generators). For each experiment models for retrained from scratch, thus due to random nature of the random walks models are significantly different.
Experiment epochs: epochs: epochs: epochs:
30+200 30+0 100+0 0+230
1 490 None 446 None
2 396 658 516 None
3 382 None None None
4 408 650 436 None
5 428 706 1046 None
6 396 None None 508
7 426 890 614 None
8 None 598 None None
9 392 464 524 None
10 None 626 452 434
Median 402 650 516 455.5
Table 2. Lengths of solutions obtained by different approaches. $S_{28}$ LRX generators. 10 experiments to decompose the ”longest” element with expected ideal length 378. Conclusion: additional training by DQN improves the results - both in length and probability to find a path. Pure DQN training is not performing well.
The analysis can be found in notebook.
3. Benchmarks
Goal - $S_{110}$. DeepMind’s success with Go game with $10^{170}$ states - creates a challenge to achieve similar results in the other fields in particular for Cayley graphs, at the moment we are somewhat far from that, but there are plenty ideas not yet incorporated in our CayleyPy project and optimistically such a goal is achievable in near horizons. Moreover in the present study we achieve pathfinding for such scale groups, but unfortunately not by fully zero prior knowledge. Our zero prior knowledge approaches achieves pathfiding for LRX generators of $S_{n}$ a bit above $30$, but at the moment $40$ is a not achieved despite bunch of neural networks we tried. What is surprising that almost all approaches we tried stuck slightly above $n=30$ all in the same region, with small variations. It is significantly better than results of computer algebra system GAP which achieves only about $n=20$, which much more lengths and timings.
3.1. One line of code to get orders of magnitude improvement from: $S_{30}$ to $S_{100}$.
if action == ’X’ and current-state[0] < current-state[1]: continue The condition used in the beam search part. Its logic is rather clear - if the contents of positions 0 and 1 are already sorted - it is logical not to swap them i.e. not to apply $X$ generator.
Surprisingly such a simple conditions is a game changer. Inserting it and using it with exactly the same pipelines which performed for around $n=30+$ we are able to find paths for $n=100$ and (apparently more - just stopped by 16G RAM limit). So the size of graphs changes from $10^{3}0+$ to $10^{1}50+$. It works the same for all models: perceptrons, gradient boostings, convolutional neural networks.
The challenge is how to avoid such prior knowledge method and extend it to other generators.
It should also be noted that this condition prevents to find optimal paths. And for small sizes of $n\leq 15$ it typically leads to a bit longer paths, but for higher values of $n$ it is a relief.
Examples of lengths found by that method are presented at table 3 below:
$S_{n}$ 100 90 80 70 60
length 16632 33459 7510 8115 3650
ideal length 4950 4005 3160 2415 1770
Time 8h 43m 3h 57m 1h 5m 52m 3h
Nb version 255 242 241 233 -
Table 3. n - permutation length. Lengths of solutions for different $n$ obtained by ML model combined followed by beam search with ”X-condition”. (Not completely zero prior knowledge approach).
Computation for $n=100$ took 8h 43m 53s GPU P100. Beam width ${2}^{20}$, 30 epochs of training. ( Notebook version 255), for other cases we used beam size not more than ${2}^{18}$, that is why for $n=90$ length is much bigger. The data for $n=60$ is obtained by the gradient boosting model, with slightly better lengths, but more timing, in the other cases single layer MLP model used. In all cases we decompose the conjecturally longest element of expected ideal length $n(n-1)/2$.
3.2. ”Overcome the GAP” with zero prior knowledge AI.
We use LRX Cayley graph of $S_{n}$ and its conjecturally longest element of expected length $n(n-1)/2$ as a benchmark. Below we demonstrate that AI methods significantly outperform classical methods of computer algebra system GAP.
$n$ GAP Ideal AI
Length Length n(n-1)/2 Length
9 41 36 36
10 51 45 45
11 65 55 55
12 78 66 66
13 99 78 78
14 111 91 91
15 268 105 105
16 2454 120 120
17 380 136 136
18 20441 153 153
19 3187 171 171
20 217944 190 190
21 - 210 210
Table 4. n - permutation length. Comparison of GAP vs AI methods (zero prior knowledge). On the conjecturally longest element of LRX Cayley graph. (Its expected ”ideal” length is $n(n-1)/2$.)
For $n=20$ GAP timing is 41min 18s, while AI methods can find results much faster. For example, simple neural network: notebook version 423: 3m 48s GPU P100. GAP benchmarks has been performed in the notebook.
$n$ Ideal AI
Length n(n-1)/2 Length
22 231 231
23 253 253
24 276 276
25 300 300
26 325 325
27 351 351
28 378 378
29 406 408
30 435 437
31 465 478
32 496 728
33 528 642
Table 5. n - permutation length. Comparison of GAP vs AI methods.
The AI results above are aggregated results from experiment with different models and approaches. In particular based on gradient boostings and neural networks, all of them quite stably and perfectly solve cases with $n\leq 30$, but going beyond that requires efforts for all our approaches.It is surprising that all types of models have the problem at the same range of $n$ and that length is not becoming gradually longer and longer, but it stays almost perfect, and then quite sharply pipeline loses ability to find a path at all. For some cases we use several reruns - if the path not found via single run, it sometimes be found by multiple relaunches of the pipeline. For example $n=32$ has been solved with our modified DQN strategy with 30 warmup epochs and 1000 DQN epochs (notebook), time: 2h 53m 53s GPU P100.
3.3. Architectures and parameter dependence.
The tests of different architectures and parameters is discussed below. The main conclusion that we observe that rather simple models like MLP and gradient boostings currently perform better, than more advanced: transformers, convolutional neural networks etc. One exception is one of our transformer models that can find paths up to $n=100$, however it is not zero prior knowledge - it uses features specific to LRX generators. Dependence on various parameters was explored in our previous paper  [Chervov25] for the case of the Rubik’s cube group, the conclusions were that beam size is one the most important parameter the length of solution is almost linearly improving on logarithm of beam size, deeper neural networks are better than wider especially, but the effect of depth is better seen on small for higher cubes, we also observed stagnation of performance with increase of training set.
For the LRX Cayley graph these phenomenons are less strongly expressed, or even negligible. Most experiments with deep neural networks have not lead to significant improvements over simple one layer perceptron with 128 neurons. Beam size effect is present, in but less expressed (see table below).
Nonbacktracking random walks have quite strong effect. In particular one can train models in just single epoch within a minute for e.g. $S_{20}$. We actually use improved version of the nonbacktracking based on the following ideas: 1) we forbid not only 1 step in history but typically 32 steps (it is a parameter) 2) we consider a bunch of trajectories and ban states for all of them at once 3) we ban not only visited states but their neighbors. The price for that is slowing down the generation, but that is not so important since gradient descent takes much longer time, and overall there is no significant slow down.
Beam Solution Length Solution Length
Width Nonbacktracking RW Simple RW
2**10 None None
2**12 252 None
2**14 202 None
2**16 194 None
2**18 194 None
2**20 190 (Ideal) None
Table 6. Without non-backtracking system cannot perform well on single epoch training. Effect of beam size is clear, but not so strong.
We generate 10000 trajectories at once each of the length 190, by the generalized nonbacktracking procedure above and train the neural network with batch size 1024 just for single epoch. Entire time for each computation - both training and beam search are 1-5 minutes, experiments can be found in versions 427-430 of the notebook.
Theoretical reason why non-backtracking is helpful is clear. Imagine our graph is a tree, then even for non-backtracking random walk number of steps coincide with the true distance (length of the shortest path). For complicated graphs it is not so, but still it more close to true distance and has less variability. I.e. the situation when the same node can be achieved in different number of steps - the fact which equivalent to having a noise in the training data and which makes models less precise.
CNN. We also analyse convolutional neural networks - however they have slower training process. They show similar results comparing to perceptrons e.g. the length of the solution for $n=16$ is 122 (ideal is 120), for $n=20$ is 216 (ideal is 190). The basic version of the architecture uses two 1D CNN layers with Batch Normalization and Dropout to avoid overfitting. One of the successful variations is using temporal gated CNN. The first part of the architecture is the residual block - several convolutional blocks with gating based on sigmoid activation function. The second part of the architecture is stacking several layers of residual blocks with gradually increased dilation to increase the size of the receptive field. Notebooks: 1, 2, 3,4.
Transformer. We also peformed experiments with a transformer model, however it was able to solve only not so big groups like $n\leq 15$. That model incorporated standard Transformer components, including multi-head self-attention, positional encoding, feed-forward layers, and dropout to mitigate overfitting. However, the model struggled to converge to a satisfactory solution, even for relatively short sequences. This issue became increasingly severe as the sequence length grew. For instance, with size = 9 (corresponding to path length = 107 and ideal = 36) or n-size = 13 (path length = 332, ideal = 78), increasing the projection dimension to 512 or 1024, adding more layers, or training for more epochs failed to address the underlying challenges. Notebooks: 1, 2.
Transformer based on LRX specific features. The features inspired by Hamming distance were used: difference between an element and its position; difference with the left neighbor; difference with the right neighbor. A standard multihead attention with one encoder layer and mlp was used to estimate the number of steps required to reach the ordered state. The transformer demonstrated the ability to handle sequences up to N=15 without using beam search. For larger N, beam search is employed, and the maximum N for which results were obtained was 100. The model and features only indirectly depend on N, therefore, when trained on micro batches constructed as random walks from an ordered permutation for N*(N-1)/2 moves, it can be used for different N values, both larger and smaller. (Notebook).
Gradient boostings. We also applied CatBoost and XGboost in our diffusion distance pipeline, surprisingly they perform not worse than neural networks (it was not the cases for Rubik’s cube - where boosting were not able to solve 4x4x4, 5x5x5 cubes), and even giving a bit shorter solutions. (Code).
3.4. Reinforcement learning on ”small graphs”
At the preliminary stage of the research we performed experiments with RL/dynamical programming methods for small graphs. In particular the question was how the initialization influence the performance. That is important to understand in view of our proposal to use initialization coming from the diffusion distance. These experiments revealed interesting phenomenons presented at figure 2. For some specific initialization the performance first degrades, and only after that achieves the convergence to the true distance which is guaranteed by the general theory. The number of steps for convergence is not more than diameter of the graph is initialization is positive, however can be larger in presence of negative values. Initialization coming from the machine learning model (”catboost”) trained on diffusion distance (same as ”warmup step” in our proposal) improves the convergence, though not in a radical manner.
Figure 2. Pearson correlation between $d_{i}(s)$ and $d(s)$ on each iteration of the DP algorithm, $\epsilon=10^{-3}$
4. Mathematical contributions
In the section we prove lower and upper bounds for the diameter and for particular elements, describe decomposition of the conjecturally longest element of the desired length $n(n-1)/2$, describe algorithms to decompose elements in LRX generators and formulate several conjectures on the growth, random walks and spectrum of the LRX Cayley graph.
4.1. LRX Cayley graph
LRX generators are: L - left cyclic shift, R - right cyclic shift, X - eXchange (flip) of positions 0,1. It can be considered as a relative of the TopSpin and Hungarian Ring puzzles (figure 3). Despite its simplicicity its structure is not fully understood. In particular according to OEIS-A186783 the diameter conjecture $n(n-1)/2$ remains open.
Figure 1 is a visualization of the LRX Cayley graph for $S_{5}$ which shows its complicated structure. More visualizations by the minimal distortion method can be found at notebook, and by originally modified force directed layout here.
Figure 3. LRX and related mechanical puzzles
4.2. Long and the longest element
Figure 4. Permutations arising from dihedral symmetries - are provebly long elements. The longest is conjecturally one of them. Surprisingly it is not the full inversion.
We created efficient implementation of the brute force breadth first search algorithm notebook using bit operations over int64 and were able to traverse all the permutations for $n\leq 14$, outperforming OEIS data by one step. From that computation we observed that the longest element is unique and have a specific form. It can be described as a symmetry of a regular polygon (figure 4), which is orthogonal to 0-1 edge. Or explicitly as product of transpositions $(0,1)(n-1,2)(n-2,3)(n-3,3)...$. We conjecture that this element is unique longest element and it has length $n(n-1)/2$. Surprisingly not the full inversion ($i\to n-i-1$) is the longest.
As we will prove below actually all the permutations which correspond to symmetries of a regular polygon are also long elements.
Proposition 4.1.
The element $(0,1)(n-1,2)(n-2,3)(n-3,3)...$ (conjecturally the longest) has the following explicit decomposition into product of LRX generators:
$\prod_{i=1}^{[(n-1)/2]}(XL^{(-1)^{(i-1)}})^{i}\prod_{i=([n/2]-1)}^{1}(L^{(-1)^{i+[(n-1)/2]-[n/2]}}X)^{i}R^{[n/2]}$ (4.1)
The proof will be given elsewhere. The decompositions of that type were first found computationally: notebook.
4.3. Lower bound on the diameter
Below, we provide a lower bound on the diameter using a combinatorial argument. The key idea is that cyclic shifts preserve the cyclic order of any given triple of numbers. Therefore, for each triple where the order is disrupted, t $X$ is required in the decomposition. The symmetries of a regular polygon, as described above, correspond precisely to the permutations that maximize changes in cyclic order. (The code to generate these long elements: notebook.)
Proposition 4.2.
Diameter of LRX graph is larger or equal $n(n-1)/2-n/2-1$
At first, we reformulate the setup as follows.
The numbers $1,\ldots,n$ are written in the vertices of a regular $n$-gon $A_{1}A_{2}\ldots A_{n}$, each number being written once. At each move, one can either rotate the arrangement by $2\pi/n$ counterclockwise or clockwise, or switch the numbers at $A_{1}$ and $A_{2}$.
For two arrangements $\pi_{1},\pi_{2}$ denote by $d(\pi_{1},\pi_{2})$ the minimal number of moves which suffice to get $\pi_{2}$ from $\pi_{1}$.
Claim. If $\pi_{2}$ is obtained from $\pi_{1}$ by an axial symmetry, then $d(\pi_{1},\pi_{2})\geqslant n^{2}/2-n-1$.
Proof. Assume that $\pi_{2}$ is obtained from $\pi_{1}$ by several moves. Call two numbers $a,b\in\{1,\ldots,n\}$ friends, if they were switched odd number of times. Since for every three numbers $a,b,c$ the orientation of a triangle formed by these numbers changed, there is odd number of pairs of friends between $a,b,c$. Let $A$ denote the set of all non-friends of element 1, $B$ the complement of $A$, i. e. $B$ consists of 1 and its friends. Our condition yields that all elements of $A$ are mutual friends, and so are all elements of $B$, but there is no friendship between $A$ and $B$. Thus, the total number of pairs of friends is $n(n-1)/2-|A|\cdot|B|\geqslant n(n-1)/2-n^{2}/4$. Thus there were at least as many switches. Between every two switches there was a rotation (otherwise we do something useless). Totally, the number of operations is not less than $n(n-1)-n^{2}/2-1=n^{2}/2-n-1$.
4.4. Upper bound on the diameter and the algorithms
We develop two algorithms which can decompose any permutation in the product of LRX generators. For the first we prove complexity bound $n(n-1)/2+3n$, thus bounding diameter from above. The code and tests for the first algorithm can be found in notebook.
Its brief description is the following:
The algorithm decomposes a permutation into the product of cycles. After that, it works with each cycle $(a_{1},a_{2},\dots,a_{n})$ in the following way:
1) We compute the permutation corresponding to the cycle.
2) We initialize the special variable $x$.
3) The element $a_{1}$ is placed in the $a_{n}$-th position, and we perform a sequence of elementary transpositions to move $a_{1}$ from the $a_{n}$-th position to the $a_{1}$-th position. During each elementary transposition, we either increase or decrease $x$ by 1. The sign of the change depends on the direction of the elementary transposition. Essentially, $x$ represents the current position of the transposition, but if our transposition is $(1,N)$, the change in $x$ is also ±$\pm 1$, even though the position changes from 1 to $N$ or from $N$ to 1.
4) Next, we attempt to restore the position of $a_{2}$ and compute the sequence of elementary transpositions that move $a_{2}$ to the $a_{2}$-th position. We apply the same operations to $x$, and so on.
5) We repeat these actions for every element of the cycle.
6) In certain situations, the elementary transposition should be replaced by $L$ or $R$ (depending on the direction in which we are moving the current element). This replacement should occur if and only if the variable $x$ satisfies the condition:
$x-c=\pm(N-1)$
Complexity estimation. First, when decomposing a cycle into transpositions, we do not shift anything by more than 1, except for elements that have already been affected by the cycle. Second, when moving an element from position $i$ (or $i\pm 1$) to position $\sigma(i)$, we perform a sequence of at most $\text{dist}(i,\sigma(i))+1$ transpositions. Third, the transition from $i\to\sigma(i)$ to $\sigma(i)\to\sigma(\sigma(i))$ takes at most one rotation. Thus, we perform at most
$\sum 2(\text{dist}(i,\sigma(i))+1)+(\text{number of elements in the cycle})$
operations. The transitions between cycles and returning back will take at most $n$ rotations. Therefore, the entire decomposition takes no more than
$2n+\sum 2(\text{dist}(i,\sigma(i)))$
operations.
Finally, after minimization,
$\sum\text{dist}(i,\sigma(i))\leq\frac{n^{2}}{4}$
which gives a total of
$2n+\frac{n^{2}}{2}+\frac{n}{2}$
operations.
We also develop a second algorithm: notebook. Currently it is our top performing algorithm from the practical tests, we expect its complexity is bounded by $n(n-1)/2+n/2$, however it is not yet proved.
It might be expected that optimal algorithm with polynomial complexity can be developed, but it is not achieved yet.
4.5. Conjectures: Gumbel for growth, spectrum uniformity, random walks mixing, etc.
Based on explicit computations for the growth for $n\leq 14$ and its analysis (notebooks 1 2 ) we come to the following conjecture, which can be thought as an analogue of the central limit theorems. It is in the vein of works by P.Diaconis et.al.  [Diaconis77, Diaconis88, Diaconis17] where for Coxeter generators growth and some other statistics were shown to follow Gaussian normal law in a limit. New point of our conjecture - appearance of asymmetric Gumbel distribution. Such asymmetry is quite natural - for the randomly chosen generators growth will be highly asymmetric with exponential growth and sharp abrupt (Rubik’s cube growth is a prototypical example), while for Coxeter generators which are nearly commutative - growth is symmetric Gaussian close to commutative case. Slight asymmetry of growth for LRX generators indicates its intermediate nature between ”exponential growth”/”commutative-like Gaussian growth” cases.
The growth - Gumbel. The growth distribution of the LRX graph follows the Gumbel distribution for large $n$.
Figure 5. Spectrum distribution of the LRX graphs.
The analysis of the spectrum of LRX graphs is performed in notebook. From the figure above it is natural to expect that spectrum tends to uniform distribution.
We have computed all possible shortest paths from the conjecturally longest element to the identity by methods of the dynamical programming (notebook ). And after that analyzed averaged trajectories (notebook) - which are analogs of random sorting networks [Virág]. The figure 6 below represents the results. It seems the pattern is different from the ”sine curves” discovered in [Virág].
Figure 6. Analogs of ”random sorting networks” for LRX. Trajectories of individual elements under average of all possible shortest trajectories between longest node and identity.
P.Diacons and L.Soloff-Coste studied random walks on LRX Cayley [Diaconis93]. They estimated that mixing time is between $n^{2}$ and $n^{3}log(n)$. We performed numerical experiments to make more precise conjecture (notebooks 1,2) unfortunately the achieved range of $n$ was not enough to draw conclusions, however it seems that mixing time is more close to $n^{3}log(n)$, moreover even for non-backtracking random walks, which is a bit surprising since they typically mix faster.
Finally we considered the Schreier coset graph for the action of $S_{2n}$ on binary strings with $n$ zeros and $n$ ones. (Analogue of $Gr(n,2n)$ over the ”field with one element”). Efficient code which stores such strings in int64 allowed us to completely traverse the graph up to $n=42$ and compute its growth. The analysis of the data presented in the notebook. Based on that it is tempting to conjecture that the diameters of such graphs are between $n^{2}/6$ and $n^{2}/5$, while growth again can be approximated by the Guimbel distribution.
5. Code availability
The code is available at the Kaggle platform, where it can be easily launched: LRX OEIS-A186783 brainstorm math conjecture
The link above leads to one of the three public Kaggle challenges which we created to stimulate the research and interplay between artificial intelligence and mathematical communities, the other two: LRX discover math and God’s algorithm, LRX-binary: in search of God’s number. Kaggle infrastructure provides a convenient way to benchmark and compare different algorithms - just making the submissions to these challenges. The code, the data and discussions are stored at the same place. Code can be executed for free on Kaggle cloud servers. The first challenge asks to decompose the conjecturally longest elements and to find their shortest decompositions - participants already achieved the length $n(n-1)/2$ an so if believe in the conjecture the optimum is found, if we do not believe - not. The second challenges asks to decompose random permutations and aims to find optimal algorithms (at the moment of writing optimum is not achieved), the third one asks the same but for the binary strings acted by LRX generators.
Acknowledgements
A.C. is deeply grateful to M. Douglas for his interest in this work, engaging discussions, and invitation to present preliminary results at the Harvard CMSA program on ”Mathematics and Machine Learning” in Fall 2024, to M. Gromov, S. Nechaev, and V. Rubtsov for their invitation to make a talk (video) at ”Representations, Probability, and Beyond: A Journey into Anatoly Vershik World” at IHES, as well as for stimulating discussions. A.C. is grateful to J. Mitchel for involving into the Kaggle Santa 2023 challenge, from which this project originated and to M.Kontsevich, Y.Soibelman, S.Gukov, A. Hayat, T. Smirnova-Nagnibeda, D.Osipov, V. Kleptsyn, G.Olshanskii, A.Ershler, J. Ellenberg, G. Williamson, A. Sutherland, Y. Fregier, P.A. Melies, I. Vlassopoulos, F.Khafizov, A.Zinovyev, H.Isambert for the discussions, interest and comments, to his wife A.Chervova and daugther K.Chervova for support, understanding and help with computational experiments.
We are deeply grateful to the many colleagues who have contributed to the CayleyPy project at various stages of its development, including: N.Bukhal, J.Naghiev, A.Lenin, E.Uryvanov, A. Abramov, M.Urakov, A.Kuchin, B.Bulatov, F.Faizullin, A.Aparnev, O.Nikitina, A.Titarenko, U.Kniaziuk, D.Naumov, A.Krasnyi, S.Botman, A.Kostin, R.Vinogradov, I.Gaiur, K.Yakovlev, V.Shitov, E.Durymanov, A.Kostin, R.Magdiev, M.Krinitskiy, P.Snopov.

### v2

@@
 Abstract.
-This paper is the second in a series of studies on developing efficient artificial intelligence-based approaches to pathfinding on extremely large graphs (e.g.
-$10^{70}$ nodes) with a focus on Cayley graphs and mathematical applications.
+This paper is the second in a series of studies on developing efficient artificial intelligence-based approaches to pathfinding on extremely large graphs (e.g., $10^{70}$ nodes) focusing on Cayley graphs and mathematical applications.
 The open-source CayleyPy project is a central component of our research.
 The present paper proposes a novel combination of a reinforcement learning approach with a more direct diffusion distance approach from the first paper.
-Our analysis includes benchmarking various choices for the key building blocks of the approach: architectures of the neural network, generators for the random walks and beam search pathfinding.
+Our analysis includes benchmarking various choices for the key building blocks of the approach: neural network architectures, random walk generators, and beam search pathfinding.
 We compared these methods against the classical computer algebra system GAP, demonstrating that they ”overcome the GAP” for the considered examples.
-As a particular mathematical application we examine the Cayley graph of the symmetric group with cyclic shift and transposition generators.
-We provide strong support for the OEIS-A186783 conjecture that the diameter is equal to n(n-1)/2 by machine learning and mathematical methods.
+We examine the Cayley graph of the symmetric group with cyclic shift and transposition generators as a particular mathematical application.
+We strongly support the OEIS-A186783 conjecture that the diameter equals n(n-1)/2 by machine learning and mathematical methods.
 We identify the conjectured longest element and generate its decomposition of the desired length.
-We prove a diameter lower bound of n(n-1)/2-n/2 and an upper bound of n(n-1)/2+ 3n by presenting the algorithm with given complexity.
-We also present several conjectures motivated by numerical experiments, including observations on the central limit phenomenon (with growth approximated by a Gumbel distribution), the uniform distribution for the spectrum of the graph, and a numerical study of sorting networks.
-To stimulate crowdsourcing activity, we create challenges on the Kaggle platform and invite contributions to improve and benchmark approaches on Cayley graph pathfinding and other tasks.
+We prove a diameter lower bound of n(n1)/2n/21n(n-1)/2-n/2-1 and an upper bound of n(n1)/2+3nn(n-1)/2+3n by presenting the algorithm with given complexity.
+We also present several conjectures motivated by numerical experiments, including observations on the central limit phenomenon (with growth approximated by the Gumbel distribution), the uniform distribution for the spectrum of the graph, and a numerical study of sorting networks.
+We create challenges on the Kaggle platform to stimulate crowdsourcing activity and invite contributions to improve and benchmark approaches to Cayley graph pathfinding and other tasks.
 Key words and phrases:
@@
 1.1. Broader context
-Deep learning has revolutionized various fields of research and was recognized by several Nobel Prizes in 2024.
+Deep learning has revolutionized various fields of research and was recognized with several Nobel Prizes in 2024.
 In recent years, machine learning has been emerging as ”a tool in theoretical science” [Douglas22], leading to several noteworthy applications to mathematical problems: [Charton19, Davies21, Bao23, Romera24, Coates23, Hayat25, Williamson24, Gukov24, Kohli25].
-The paper presents advancements in applying artificial intelligence methods to the mathematical problem of Cayley graph pathfinding, which is equivalent to decomposing a group element into a product of generators.
-It is the second work in a series devoted to the CayleyPy project to develop efficient machine learning-based approaches and an open source library to deal with extra large graphs (e.g.
+The paper presents advancements in applying artificial intelligence methods to the mathematical problem of Cayley graph pathfinding, equivalent to decomposing a group element into a product of generators.
+It is the second work in a series devoted to the CayleyPy project to develop efficient machine learning-based approaches and an open-source library to deal with extra-large graphs.
+(e.g.
 $10^{70}$) with a focus on finite Cayley graphs and mathematical applications.
@@
 They serve as both a prototype and an inspiration for many developments, including ours.
-In a nutshell, the method consists of two main components: a neural network trained to suggest moves from any given state, and a graph pathfinding algorithm that helps to find a path even when the neural network’s predictions are not sufficiently precise.
+The method consists of two main components: a neural network trained to suggest moves from any given state and a graph pathfinding algorithm that helps to find a path even when the neural network’s predictions are not sufficiently precise.
 The core idea is similar across most approaches, but achieving optimal performance for each task requires analyzing and benchmarking various neural networks and graph pathfinding methods.
-One of the goals of this work is to present such an analysis for some particular Cayley graphs.
-1.2. Present paper: approach, benchmarks, mathematical theorems and conjectures
+Graph Group RL
+vertex / node permutation / element state
+edge generator action
+path word trajectory
+shortest path shortest word maximal return
+diameter longest element ?
+Table 1. Comparison of graph theory, algebraic and reinforcement learning terminology.
+One of the goals of this work is to apply these methods to a particular class of Cayley graphs.
+As in our previous paper [Chervov25], we choose a set of generators $a_{1}^{\pm 1},\ldots,a_{k}^{\pm 1}\in S_{n}$.
+All finite products (words) of these elements form a subgroup $G\subset S_{n}$ which we identify with its Cayley graph.
+Due to vertex-transitivity of Cayley graphs, the pathfinding problem can be reduced to searching a path from some element $g\in G$ to the identity permutation $e$; this is the same as finding a decomposition of $g$ into a word $\prod\limits_{j=1}^{m}a_{i_{j}}^{\pm 1}$, $1\leq i_{j}\leq k$.
+The problem of finding the shortest path (or word) can also be reformulated in terms of a generic reinforcement learning approach discussed in Section 2.1.
+A set of dictionary between RL, graphs, and group is presented in Table 1.
+1.2. Present case study - LRX Cayley graph.
 Figure 1. LRX Cayley graph for $S_{5}$
-The paper presents both AI contributions and mathematical results and conjectures.
-Here, we work with a particular Cayley graph defined by the L - left cyclic shift, R - right cyclic shift and X - eXchange of positions $0,1$, it will be denoted LRX following OEIS-A186783.
-Its image for $S_{5}$ is shown on figure 1.
-Developments of AI methodology.
-We propose a modified deep Q-learning (DQN) method for the case of the graph pathfinding task which resolves the so-called sparse reward problem of the standard DQN approach for graph pathfinding.
-The proposal combines DQN with the more simple diffusion distance approach from our previous paper  [Chervov25].
-An analysis and benchmarks of these approaches are presented, including tests of various neural network architectures such as transformers, convolutional networks, and perceptrons.
-It is demonstrated that these methods allow finding paths for $S_{n}$ up to $n$ around 30 - for the particular class of generators (LRX) which are in the main focus of the present paper.
-We demonstrate that they ”overcome the GAP” - outperform classical computer algebra system GAP on the same task which can work only up to $n$ around 20.
+The paper presents both AI contributions and mathematical results and conjectures. We use LRX Cayley graphs as primary benchmark in the subsequent discussion.
+LRX Cayley graphs. For each $n\in\mathbb{N}$ the symmetric group $S_{n}$ can be generated by three permutations:
+  1. (1)
+    left cyclic shift $L=(1\ 2\ 3\ \ldots\ n-1\ 0)$;
+  2. (2)
+    right cyclic shift $R=L^{-1}=(n-1\ 0\ 1\ 2\ \ldots\ n-3\ n-2)$;
+  3. (3)
+    transposition (eXchange) of the first two positions $X=X^{-1}=(0,1)$.
+These generators and the corresponding Cayley graph will be denoted LRX following OEIS-A186783.
+Visualization (Fig.
+1) of the LRX Cayley graph for $S_{5}$ reveals its complicated structure.
+LRX Cayley graph can be considered as a relative of the Oval Track (or TopSpin) and Hungarian Rings puzzles (Fig.
+2).
+Figure 2. LRX and related mechanical puzzles
+Conjectures on diameters of LRX and LX (V.M.
+Glushkov’s 1968 problem).
+It is conjectured OEIS-A186783 that diameters of LRX are n(n1)/2n(n-1)/2, that conjecture will be in the focus of the present paper.
+The case of two generators just and L and X (not inverse closed) goes back at least to 1968 influential paper by “one of the founding fathers of Soviet cybernetics” V.
+M.
+Glushkov [Glushkov68], which has been much studied (survey: [GlukhovZubov99] pages 18-21), but not resolved.
+The conjecture on the diameter in that case is presented in our subsequent paper: [CayleyPyGrowth].
+Not exactly in the present scope, but worth mentioning: there was a long-standing problem resolved positively recently [SawadaWilliams19] of whether the directed graph generated by L and X (also denoted Sigma and Tau) has a hamiltonian path.
+D.Knuth assigned to that problem complexity 48 out of 50, it was first articulated in [NijenhuisWilf75].
+1.3. AI contributions.
+Development of AI methodology.
+We suggest a modified deep Q-learning (DQN) method for the case of the graph pathfinding task, which resolves the so-called sparse reward problem of the standard DQN approach for graph pathfinding.
+The proposal combines the DQN and the diffusion distance approach from [Chervov25] and presents an analysis and benchmarks of these techniques, including tests of various neural network architectures such as MLP, CNN, and transformers.
+These methods have been shown to find paths for LRX Cayley graphs up to $n$ around 40 and with small modification (X-trick, which is specfic to LRX) even beyond 100.
+We demonstrate that they ”overcome the GAP”: outperform the classical computer algebra system GAP on the same task, which can work only up to $n$ around 20, also producing much more optimal paths and working faster.
 Efficient PyTorch implementation.
-We pay significant attention to develop an original highly optimized code, the current version is specific to Cayley graphs of permutation groups (matrix groups will be supported later).
+We devote significant attention to developing an original, highly optimized code.
+The current version is specific to Cayley graphs of permutation groups (matrix groups will be supported later).
 We propose many technical solutions for fast parallel processing of large amounts of permutations, including product computations, efficient hashing, and extracting unique states.
 The code works on both CPU and GPU without modification.
-One line of code bringing orders of magnitudes improvement.
-We report a curious finding - a single-line modification in the code of our main graph pathfinding module (beam search), dramatically improves performance.
-This enhancement increases the feasible pathfinding size from symmetric groups $S_{n}$ with $n$ to $n$ around 100 (and apparently even further with more hardware).
+X-trick.
+We report a curious finding: a single-line modification (which we call X-trick) in the code of our main graph pathfinding module (beam search) dramatically improves performance.
+This enhancement increases the feasible pathfinding size to $n$ around 100 (apparently even further with more powerful hardware).
 Unfortunately, this improvement is currently specific to the particular LRX Cayley graphs considered in this paper, but we hope it can be generalized.
-Mathematical contributions.
-LRX Cayley graphs.
-Diameter conjecture OEIS-A186783.
-In the present paper, we focus on a specific family of Cayley graphs for the symmetric groups $S_{n}$ those generated by the two cyclic shifts (L) - left, (R) - right and the transposition X=$(0,1)$ of the first two positions.
-Surprisingly, according to OEIS-A186783, it is an open conjecture that the diameter of the Cayley graph is $n(n-1)/2$ for $S_{n}$.
-These generators and Cayley graphs will be denoted LRX as in OEIS.
-The longest element.
-By implementing an efficient brute-force breadth-first search algorithm, we traverse the full LRX graphs for $n\leq 14$, confirming that the longest element has a length of $n(n-1)/2$, as predicted.
-Moreover, we observe that this element is always unique and follows a clear pattern - it is a product of transpositions: $(0,1)(n-1,2)(n-2,3)(n-3,4)...$, so we expect this for all $n$.
+Moreover, we provide a mathematical counterpart of that finding - X-trick modified random walks are mixing much faster - see discussion below.
+1.4. Mathematical contributions.
+We present several conjectures about LRX Cayley graphs which could leverage a more general understanding of the properties of that Cayley graph and its relatives.
+Diameter conjecture and the longest element.
+Surprisingly, according to OEIS-A186783, it is an open conjecture that the diameter of the Cayley graph is n(n1)/2n(n-1)/2 for $S_{n}$.
+By implementing an efficient brute-force breadth-first search algorithm, we traverse the full LRX graphs for $n\leq 15$, confirming that the longest element has a length of n(n1)/2n(n-1)/2, as predicted.
+Moreover, we observe that this element $\ell_{n}$ is always unique and follows a clear pattern
+$\ell_{n}=(1\ 0\ n-1\ n-2\ \ldots\ 3\ 2)=(0,1)(n-1,2)(n-2,3)\ldots=R^{2}e^{\leftarrow},$ (1)
+where $e^{\leftarrow}=(n-1\ n-2\ \ldots\ 2\ 1\ 0)$ is reversed identity permutation.
+We expect (1) holds for all $n$.
+Interestingly, the longest element $\ell_{n}$ is equal not to the reversed identity permutation itself but to the element obtained by two right shifts from it.
 AI and crowd-sourcing to support the conjecture.
-For that particular element, we have launched various versions of our pipelines and have consistently found that the decomposition is never shorter than $n(n-1)/2$.
-We organized a challenge on the Kaggle platform giving that particular element for $n=2...200$ and requesting decompositions of the shortest possible length.
-Once again, no participant found a decomposition shorter than $n(n-1)/2$, thereby confirming the expectation.
-Rigorous lower and upper bounds.
-We prove lower bound $n(n-1)/2-n/2-1$ by a combinatorial argument (improving $n(n-1)/3$ [Babai89]); afterwards, we prove the diameter upper bound $n(n-1)/2+3n$ by the algorithm with such complexity (improving $3/2n^{2}$  [Kuppili20]).
-We also develop another algorithm which empirically shows better complexity: $n(n-1)/2+n/2$, but lacks the rigorous proof for such an estimate.
-Finally, we present explicit decomposition for the conjecturally longest element of the desired length $n(n-1)/2$.
-Central limit phenomena for growth - Gumbel distribution.
-Computations suggest the following conjecture: for large $n$, the growth function of LRX graphs follows an asymmetric Gumbel distribution.
-This is in the vein of field-shaping works by P.
+For the longest element $\ell_{n}$, we have launched various versions of our pipelines, and have consistently found that a decomposition is never shorter than n(n1)/2n(n-1)/2.
+We organized a challenge on the Kaggle platform suggesting to decompose $\ell_{n}$, $n=2,\ldots,200$, into the shortest possible word.
+Once again, no participant found a decomposition shorter than n(n1)/2n(n-1)/2, thereby confirming the expectation.
+Rigorous lower and upper bounds for the diameter.
+We prove the lower bound n(n1)/2n/21n(n-1)/2-n/2-1 by a combinatorial argument (improving n(n1)/3n(n-1)/3 [Babai89]); later, we prove the upper bound n(n1)/2+3nn(n-1)/2+3n by the algorithm with such complexity (improving $\frac{3}{2}n^{2}$ [Kuppili20]).
+We also develop another algorithm that empirically shows better complexity n(n1)/2+n/2n(n-1)/2+n/2, but it lacks a rigorous proof for such an estimate.
+Finally, we present an explicit decomposition for the conjecturally longest element $\ell_{n}$ of the desired length n(n1)/2n(n-1)/2.
+Central limit phenomena for growth.
+Computations suggest the following conjecture: for large $n$, the growth function of LRX graphs follows a left-skewed Gumbel distribution.
+This statement is in the vein of field-shaping works by P.
 Diaconis et.
 al.
-[Diaconis77, Diaconis88, Diaconis17], who demonstrated that the growth of the neighbor transposition graph and many related statistics asymptotically follow the Gaussian normal distribution.
-We study numerically the spectrum of LRX graph and observe surprisingly uniform distribution of eigenvalues, we also analyze statistics of the paths from the longest element to the identity (”sorting networks”) we observe numerically that pattern differs from the famous ”sine curves” for Cayley graph of neighbor transpositions [Virág]; numerical analysis of random walks shows consistency with results of Diaconis [Diaconis93] on mixing time for LRX Cayley graph.
-We also present similar analysis for the Schreier coset graph for the action of the LRX generators on binary strings consisting of $n/2$ zeros and ones (for even $n$).
+[Diaconis77, Diaconis88, Diaconis17], who demonstrated that the growth of the neighbor transposition graph and many related statistics asymptotically follows the Gaussian normal distribution.
+However we observe that skewness on the growth distribution for LRX does not tend to zero, thus Gaussian is not a choice, while Gumbel provides a good numerical fit.
+We also present a similar analysis for the Schreier coset graph for the action of the LRX generators on binary strings consisting of $n/2$ zeros and ones (for significantly larger range of $n$).
+Random walks and mixing time.
+P.Diaconis and L.Saloff-Coste [Diaconis93] obtained an important result that mixing time for LRX is between $O(n^{3})$ and $O(n^{3}log(n))$.
+We support their result numerically and complement it by considering two other types of random walks: non-backtracking and X-trick modified.
+Both show faster mixing, however non-backtracking is only faster by a constant factor, while X-trick leads to around $O(n^{2})$ mixing.
+Which particially explains its effectivity in beam search.
+Spectrum and sorting networks.
+We study numerically the spectrum of the LRX graph and observe a surprisingly uniform distribution of eigenvalues.
+Investigating statistics of the paths from $\ell_{n}$ to $e$ (”sorting networks”), we observe numerically that the pattern differs from the famous ”sine curves” for Cayley graph of neighbor transpositions [Virág].
 In summary, our analysis demonstrates that computational experiments, especially those involving AI, can be very helpful in advancing mathematical research on Cayley graphs.
-AI methods significantly outperform classical methods of computer algebra systems such as GAP, performing well for $S_{n}$ with $n$ around 30, and with a small addition of prior knowledge up to 90, while classical methods perform well only up to 20.
-1.3. Finite Cayley graphs: pathfinding techniques, results and open problems
-Let us briefly summarize existing techniques for Cayley graphs pathfinding, and remind more general context on Cayley graphs research.
+AI methods significantly outperform classical methods of computer algebra systems such as GAP, performing well for $S_{n}$ with $n$ around 40 and with a small addition of prior knowledge up to 90, while classical methods perform well only up to 20.
+1.5. Review. Finite Cayley graphs: pathfinding techniques, results and open problems
+Let us briefly summarize existing techniques for Cayley graphs pathfinding, and provide more general context on Cayley graphs research.
 Optimal pathfinding is typically NP-hard.
 Finding the shortest paths on generic finite Cayley graphs is an NP-hard problem [Goldreich81] (even P-space complete [Jerrum85]).
-As is the case for many specific group families, such as NxNxN Rubik’s Cube groups [Demaine17] and others [Bulteau15].
-For particular groups it is hard to develop practical optimal algorithm e.g.
-standard 3x3x3 Rubik’s cube the first optimal algorithm has been proposed only 1997 [Korf97].
-It required 192 Gigabytes of precomputed data and 12-24 hours to solve single cube (methods were later improved  [^1] M.Reid: https://www.cflmath.com/Rubik/optimal_solver.html, H.Kociemba 2021: https://pypi.org/project/RubikOptimal/, Andrew Skalski: https://github.com/Voltara/vcubem , [Rokicki14] to several cubes per second).
-To develop practical optimal solver for 4x4x4 Rubik’s cube - is a challenge.
+As is the case for many specific group families, such as $N\times N\times N$ Rubik’s Cube groups [Demaine17] and others [Bulteau15].
+It is hard to develop a practical optimal algorithm for particular groups, for instance, the first optimal algorithm for the standard $3\times 3\times 3$ Rubik’s cube has been proposed only in 1997 [Korf97].
+It required 82 Megabytes of precomputed data and 12-24 hours to solve a single cube (later improved  [^1] M.Reid: https://www.cflmath.com/Rubik/optimal_solver.html, H.Kociemba 2021: https://pypi.org/project/RubikOptimal/, Andrew Skalski: https://github.com/Voltara/vcubem , [Rokicki14] to several cubes per second).
+Developing a practical optimal solvers for a $4\times 4\times 4$ Rubik’s cube is challenging open problem.
 Non-optimal pathfinding: Schreier-Sims algorithm, GAP implementation, non-effectivity for large groups.
-The Schreier-Sims algorithm [Sims70],[Knuth91] (or its randomized versions) is the method typically used for decomposing group element into product of generators for permutation groups.
-It is implemented in GAP computer algebra system.
-However its outputs ”are usually exponentially long” [Shamir89] and that forbids practical computations for large groups (typically the sizes beyond $10^{30}-10^{40}$ are out of reach - examples and analysis is given below).
+The Schreier-Sims algorithm [Sims70] (and its randomized version [Knuth91]) is the method typically used for decomposing group elements into products of generators for permutation groups.
+It is implemented in the GAP computer algebra system.
+However, its outputs ”are usually exponentially long” [Shamir89], and that forbids practical computations for large groups (typically the sizes beyond $10^{30}$$10^{40}$ are out of reach – examples and analysis are given below).
 Non-optimal pathfinding: ”small support” Kaggle Santa 2023 methods (generality issue).
-The methods which can deal with extremely large groups (e.g.
-size $>10^{1000}$ like 33x33x33 Rubik’s cube) were demonstrated by top participants of the recent challenge organized on the Kaggle platform - Santa 2023.
-However the generality of that approach is unclear.That is contrast to AI methodology which is from the beginning is general not even restricted to Cayley graphs.
+The methods which can deal with extremely large groups (e.g., $10^{1000}$ in size like $33\times 33\times 33$ Rubik’s cube) were demonstrated by top participants of the recent Kaggle challenge Santa 2023.
+However, the generality of that approach is unclear.
+That is in contrast to AI methodology, which is, from the beginning general not even restricted to Cayley graphs.
 The idea of ”Santa methods” is quite noteworthy: one should first find ”small support elements” achievable from original generators and take them as new generators, then one simply runs beam search with Hamming distance heuristics to find path.
-Method becomes effective since for small support generators Hamming distance and true graph distance are quite related by obvious reason.
-”Small support” are those permutations which change only small number of positions.
-They are well-known for Rubik’s cube solvers - ”commutator” or ”Pif-Paf” moves [Mulholland16], M.Brandenburg post.
-In mathematics small support elements play an important role in deep works by Babai, Seress, Helfgott, et.al.
-[Babai88, Babai04, Seress14, Helfgott14, Helfgott19] (slides: A.Seress,H.Helfgott ), however it is unclear whether one can effectively find small support elements from given generators, thus making unclear the generality of ”small support methods”.
+This method becomes effective since for small support generators, Hamming distance, and true graph distance is quite related for obvious reasons.
+”Small support” is those permutations that change only a small number of positions.
+They are well known for Rubik’s cube solvers – ”commutator” or ”Pif-Paf” moves [Mulholland16], M.Brandenburg post.
+In mathematics, small support elements play an important role in deep works by Babai, Seress, Helfgott, et.al.
+[Babai88, Babai04, Seress14, Helfgott14, Helfgott19] (slides: A.Seress,H.Helfgott).
+However, the question whether one can effectively find small support elements from given generators remains open, thus making the generality of ”small support methods” unclear.
 Non-optimal pathfinding: algorithms for specific families of generators.
-For some particular choices of generators there are algorithmic solvers which can effectively find non-optimal solution.
+Some algorithmic solvers can effectively find non-optimal solutions for some particular generator choices.
 It seems unlikely that one can produce similar algorithms for any generators.
-The prototypical example is bubble sort algorithm, which solves the problem for the neighbor transpositions generators of $S_{n}$.
-Algorithmic solvers exist for Rubik’s cube of arbitrary sizes and for other puzzles related groups.
-The other families of generators and algorithms are important in bioinformatics for estimation of the evolutionary distance.
-These generators are related to flips of subsequences - see e.g.
-[Pevzner95, Pevzner99, Bulteau19].
-Related example is pancake sorting (or prefix sorting) problem with the first algorithm proposed by Bill Gates et.al.
+The prototypical example is the bubble sort algorithm, which solves the problem for the neighbor transposition generators of $S_{n}$.
+Algorithmic solvers exist for Rubik’s cube of arbitrary sizes and other puzzle-related groups.
+The other families of generators and algorithms are important in bioinformatics for estimating the evolutionary distance.
+These generators are related to flips of subsequences (see, e.g., [Pevzner95, Pevzner99, Bulteau19]).
+A related example is pancake sorting (or prefix sorting) problem with the first algorithm proposed by Bill Gates et.al.
 [Gates79].
 AI-methods for pathfinding.
-To the best of our knowledge there is no systematic effort for Cayley graph pathfinding with AI methods except our CayleyPy project.
-However for some particular cases for example Rubik’s cube there exists a number of approaches, e.g.
-based on genetic algorithms: [Świta23].The most notable are: DeepCube [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] - the first AI based successful solver for 3x3x3 cube (website), the second one: EfficientCube [Takano21], some other: [TakanBrunetto17, Johnson21, Amrutha22, Noever21, Chasmai21, Bedaywi23, Pan21] which not achieve complete solution.
-Noteworthy idea [Pan21] is to combine neural networks with representations theory of the symmetric group — neural net predicts the coefficients of the non-abelian Fourier transform for the distance function.
+To the best of our knowledge, there is no systematic effort for Cayley graph pathfinding with AI methods except for our CayleyPy project.
+However, for some particular cases, for example, Rubik’s cube, there are several approaches, e.g., based on genetic algorithms: [Świta23].The most notable are: DeepCube [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] — the first AI based successful solver for $3\times 3\times 3$ cube (website), the second one: EfficientCube [Takano21], some other: [TakanBrunetto17, Johnson21, Amrutha22, Noever21, Chasmai21, Bedaywi23, Pan21] which not achieve complete solution.
+Noteworthy idea [Pan21] is to combine neural networks with the representation theory of the symmetric group — neural net predicts the coefficients of the non-abelian Fourier transform for the distance function.
 The rationale is: observed sparsity (bandlimitedness) of Fourier transform of the common distance functions on $S_{n}$  [Swan17].
-The recent seminal achievement from S.Gukov’s team:  [Gukov24] creates effective AI based pathfiding approach for infinite group of Andrews-Curtis moves and resolves the Akbulut-Kirby conjecture remained opened for 39 years.
-AI-methods are also used for pathfinding in the context of planning movements in obstacle-rich environments and road networks it is related field, but with different focus [Leskovec22, Yakovlev23]
+The recent seminal achievement from S.Gukov’s team:  [Gukov24] creates an effective AI-based pathfinding approach for an infinite group of Andrews-Curtis moves and resolves the Akbulut-Kirby conjecture remained open for 39 years.
+AI methods are also used for pathfinding in the context of planning movements in obstacle-rich environments and road networks  [Leskovec22, Yakovlev23].
+After the first version of the present work, several new contributions appeared: [Douglas25] proposed a novel diffussion-like approach to that problem, considered several classes of permutation and matrix groups and also demonstrated that the deep learning approach works successfully not only on Rubik’s cube, but on general Cayley graphs up to certain sizes.
+[Ziarko25] provided solution of 333 Rubik’s cube based on approach of latent representations.
 Cayley graphs: applications, diameters, random walks, open conjectures.
-Let us give a more general context on Cayley graph research.
-Cayley graphs are fundamental in group theory [Gromov93],[Tao15], and have various applications: bioinformatics  [Pevzner95, Pevzner99, Bulteau19]; processors interconnection networks [Akers89, Cooperman91, Heydemann97];in coding theory coding theory and cryptography [Wigderson06, Zémor94, Petit11]; in quantum computing [Kohli24, Sarkar24, Vidick23, Acevedo06, Gromada22], etc.
-Two focuses of pure mathematical research on finite Cayley graphs are related to diameters estimation and the behavior of random walks.
-In the case of the symmetric group $S_{n}$ there are two open conjectures which are easy to formulate, but somewhat representative for the field.
-Babai-like conjecture: diameter of $S_{n}$ is not more than $O(n^{2})$ for any choices of generators (see e.g. H.Helfgott’s surveys:  [Helfgott14],[Helfgott19],[Helfgott15] ;
-Diaconis  [Diaconis13] conjecture: the mixing time for random walks is not more than $O(n^{3}log(n))$ (again for any choices of generators).
-The first one can be thought as a particular case of the Babai conjecture [Babai92], [Tao15] which predicts that diameter of simple groups is not as large as one may expect: $O(log^{c}|G|)$ for some absolute $c$ — in contrast to abelian groups where it can be $O(|G|)$.
-Thus conjecture represents an intriguing connection between algebraic properties of a group (which does not depend on the choice of generators) - being simple and the property of Cayley graphs - having short diameter (which is generator dependent property).
+Let us provide more general context on Cayley graph research.
+Cayley graphs are fundamental in group theory [Gromov93],[Tao15], and have various applications: bioinformatics  [Pevzner95, Pevzner99, Bulteau19]; processors interconnection networks [Akers89, Cooperman91, Heydemann97]; in coding theory and cryptography [Wigderson06, Zémor94, Petit11]; in quantum computing [Kohli24, Sarkar24, Vidick23, Acevedo06, Gromada22], etc.
+Pure mathematical research on finite Cayley graphs focuses on diameter estimation and the behavior of random walks.
+In the case of the symmetric group $S_{n}$, two open conjectures are easy to formulate, but somewhat representative of the field:
+  • Babai-like conjecture: for any choices of generators diameter of $S_{n}$ is $O(n^{2})$ (see, e.g.,  [Helfgott14],[Helfgott19],[Helfgott15]);
+  • Diaconis  [Diaconis13] conjecture: the mixing time for random walks is $O(n^{3}\log n)$ (again for any choices of generators).
+The first one can be thought of as a particular case of the Babai conjecture [Babai92], [Tao15], which predicts that diameter of simple groups is not as large as one may expect: $O(\log^{c}|G|)$ for some absolute $c$ — in contrast to abelian groups where it can be $O(|G|)$.
+This conjecture represents an intriguing connection between an algebraic property of a group — simplicity, which is independent of the choice of generators — and a property of its Cayley graphs — having small diameter, which does depend on the generators.
 Estimating diameters of Cayley graphs is a hard problem.
-One of its applications is latency estimation for processor interconnection problem [Akers89],[Cooperman91], [Heydemann97] - since diameter corresponds to maximal latency.
-It is also of interest for community of puzzles enthusiasts - like Rubik’s cube.
-Diameter is also called a ”God’s number” - the number of moves the worst configuration can be solved by the best (optimal) algorithm.
-It is widely unknown for most of the puzzles, its determination for the Rubik’s cube 3x3x3 for the standard choices of generators required significant efforts and computational resources [Rokicki14].
-It is not known precisely for higher cubes, neither for 3x3x3 Rubik’s with STM, ATM metrics.A precise conjectures on 4x4x4 and 5x5x5 Rubik’s cube diameters appeared recently [Hirata24], collections of conjectures and estimates for many puzzles can be found on speedsolving website.
-As we will discuss below novel AI methods might provide insights via computational experiments to such kind of questions.
-2. Reinforcement learning (RL) and graphs pathfiding
-2.1. Reminders: RL, sparse reward problem, Bellman equation for graph distances, Q-algorithm, pathfinding
-RL in brief.
-Reinforcement learning is an important direction in AI that differs from supervised and unsupervised learning in the following ways: it focuses on optimizing cumulative rewards, and it requires the agent to, in a sense, ”create its own training set”.
-The agent must choose the best strategies to ”explore the space” and select training examples on its own, in contrast to supervised learning, where the training set is provided in advance.
+One of its applications is latency estimation for the processor interconnection problem [Akers89],[Cooperman91], [Heydemann97] since diameter corresponds to maximal latency.
+It is also a question of interest for community of puzzles enthusiasts — like Rubik’s cube.
+In the context of puzzles the diameter is often called the ”God’s number” — the number of moves the best algorithm can solve the worst configuration.
+It is widely unknown for most of the puzzles.
+Its determination for the Rubik’s cube $3\times 3\times 3$ for the standard choices of generators required significant efforts and computational resources [Rokicki14].
+It is not known precisely for higher cubes, neither for $3\times 3\times 3$ Rubik’s with STM, ATM metrics.
+Precise conjectures on $4\times 4\times 4$ and $5\times 5\times 5$ Rubik’s cube diameters appeared recently [Hirata24], collections of conjectures and estimates for many puzzles can be found on speedsolving website.
+As we will discuss below, novel AI methods might provide insights into such questions through computational experiments.
+2. Reinforcement learning (RL) and graphs pathfinding
+The last subsection (2.4) presents our new proposal for combining reinforcement learning (RL) training with pretraining based on the diffusion distance approach introduced in our previous work [Chervov25], which is briefly reviewed in Section 2.2.
+Both methods rely on beam search, recalled in Section 2.3.
+The first subsection (2.1) provides a brief overview of reinforcement learning and explains how it naturally fits into the graph pathfinding framework — a simple yet not thoroughly discussed perspective in the existing literature, to the best of our knowledge.
+2.1. RL setup for pathfinding and Bellman equation
+Reinforcement learning is an important subfield of AI, which is different from supervised and unsupervised learning in the following respects: it aims to optimize cumulative rewards, and it requires the agent to, in a sense, ”create its own training set”.
+The agent should choose the best strategies to ”explore the space” and select training examples by himself, in contrast to supervised learning, where the training set is given a priori.
+Here we will briefly discuss it from the graph theory perspective.
+Bellman equation and shortest paths.
+Let $d(s)$ be the distance function — the length of the shortest path from the state $s$ to the target node $e$.
+Denote the set of all neighbors of $s$ as $N(s)$.
+It is obvious that, the distance function $d(s)$ satisfies the Bellman equation:
+$d(s)=1+\min\limits_{t\in N(s)}d(t)$ (2)
+and the boundary condition
+$d(e)=0.$ (3)
+The equation is elementary but plays an important role in reinforcement learning and its deep learning generalizations discussed below.
+The Bellman equation implies a simple proposition, which is easy to see inductively considering neighbor nodes to $e$, then neighbors of neighbors, etc.
+(actually, this is how the breadth-first search algorithm finds the shortest path).
+Proposition 2.1.
+For any finite connected graph $G$, the only solution to the equations (2) and (3) is given by the distance function $d(s)$.
+Moreover, the only solutions to Bellman’s equation (2) (without boundary condition) correspond to the distance to some sets of nodes in $G$ (up to adding a constant).
+That set can be recovered as minimums of $d(s)$.
 RL ”equivalence” to graph pathfinding.
-The setup of reinforcement learning is very similar to the setup of graph pathfinding: states - are nodes of the graph, actions corresponds to edges/transitions to neighbor states by these edges, rewards corresponds to weights on the edges.
-The cumulative reward corresponds to length of paths (with respect to edges weights).
-Typically in graph theory shortest paths are of interest - that means ”reward” is minimized, not maximized as in RL, so it is better to use term not the ”reward”, but ”penalty”.
-Unweighted graphs and sparse reward problem.
-The sparse reward problem is a well-known challenge in reinforcement learning (RL), and it specifically arises in pathfinding on Cayley graphs and, more generally, on any unweighted graph, which makes standard RL approaches perform poorly.
-The sparse reward problem occurs when rewards are given very rarely, making it difficult to generate data for training.
-This situation is particularly prevalent in graph pathfinding on large graphs, where the task is to find a path to a single node out of trillions of others.
-A reward is only given once the desired node is reached, but it is practically impossible to find it with a ”cold start” through random exploration.
-It’s important to note that the condition of all weights being equal to one plays a significant role.
-If the weights were different, one could reasonably adopt a strategy of selecting the step with the best local reward.
-However, when all weights are identical, there is no local preference to guide the agent in making the right move.
-There are several strategies to address this issue, and below, we propose a new approach.
-Bellman value equation is equation for distance functions.
-The value function (or ”position estimation”) in RL exactly corresponds to the distance functions on graphs.
-To make that analogy more precise: consider an unweighted graph $G$ (all edges weights are equal to 1), choose some node $e$ (e.g.
-identity of a group in Cayley graph), consider a function $d(g)$ the distance of the shortest path from $g$ to $e$ (it is ”word metric” in group theory).
-Then there is a simple relation:
-$d(g)$ $=1+\min_{n:~{}neigbours~{}of~{}g}d(n)\text{~{}~{}~{}~{}~{}~{}~{}~{}~{} Bellman equation}$ (2.1)
-$d(e)$ $=0\text{~{}~{}~{}boundary condition}$ (2.2)
-This relation is simple, but it is fundamental - it is a particular case of the Bellman’s equation.
-The properties can be summarized in the following simple proposition, which is easy to see inductively considering neghbor nodes to $e$, then neighbors of neighbors and so on.
-Proposition 2.1.
-For any finite connected graph $G$ the only solution to the pair of equations above are given by the distance function (length of the shortest path) to the node $e$.
-Moreover the only solutions to the Bellman’s equation (without boundary condition) correspond to distance to some sets of nodes in $G$ (up to adding a constant).
-That set can be recovered as minimums of $d(g)$.
-Sense of Bellman equation - local condition for globally (cumulatively) defined object.
-The Bellman equation is equation for the length of the shortest path - that is globally/cumulatively defined object, but the condition itself is local - only at the neighbor of each point.
-Idea of Q-learning and deep Q-learning algorithms for value/distance function.
-These algorithms provide a way to compute solutions of the Bellman equation, which means to find the distance functions on graphs.
-One starts with some initialization for $d(g)$, then compute next iteration:
-$d_{k}(g)=1+\min_{n:~{}neigbours~{}of~{}g}d_{k-1}(n),d_{k}(e)=0$ by applying the right hand side of the equations 2.1.
-There are theoretical guarantees that such a process will converge to the solution of the Bellman equation (and boundary condition).
-Idea of deep Q-learning algorithm.
-The deep Q-learning is used in the situations where the graph is large (can be of google size).
-One assumes that for each node of graph there is a feature vector attached.
-One considers a neural network which takes a feature vector as an input and should output the approximate solution for Bellman equations.
-The training process is organized as follows.
-Each epoch one selects some subset of nodes and from current neural network $d(g)$ computes the right hand side of the equations 2.1: $t(g)=1+\min_{n:~{}neigbours~{}of~{}g}d(n),t(e)=0$, denoting $t(g)$ as a result of that computation.
-Afterwards one runs gradient descent to minimize the error between $d(g)$ and $t(g)$.
-If the error would be zero that means that $d(g)$ solves the Bellman equations.
-And thus in graph theory setup it produces the distance function (”word metric” in the case of Cayley graphs).
-Typically the choice of the training subset of nodes is made with the help of walking over the graph combining both random steps and steps guided by the model.
-We will discuss that in details below.
+Here we remind the standard form of Bellman equation for the value function $v_{*}(s)$ ([Sutton18, p.62]) and relate it to the form given above.
+Taking any action $a$ in state $s$ the agent receives a reward (penalty) $r(s,a)$, and denote discount factor to be $\gamma\in(0,1]$, then Bellman optimality equation:
+$v_{*}(s)=\max\limits_{a}\big\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}}v_{*}(s^{\prime})\big\}$
+Considering environment to be deterministic, $\mathbb{E}_{s^{\prime}}v_{*}(s^{\prime})=v_{*}(a\circ s)$ and taking r(s,a)=1r(s,a)=-1, one gets:
+$v_{*}(s)=-1+\gamma\max\limits_{a}\big\{v_{*}(a\circ s)\big\}.$ (4)
+Now observe that (2) turns into (4) if $\gamma=1$ and $d(s)=-v_{*}(s)$. By definition of the value function $v_{*}(e)=0$ which is the same as (3).
+Solving Bellman - simple iteration method (dynamic programming).
+Bellman equation on unknown function $d$ has the form $d=F(d)$, thus to solve it one can use method of simple iteration: $d_{k+1}=F(d_{k})$.
+Thus, solving Bellman equation is conceptually rather simple - starting with some initialization $d_{0}(s)$, the new approximation for the distance function is iteratively updated by by the right hand side of the Bellman equation (method is also called dynamic programming in RL context), i.e.:
+$d_{k}(s)=1+\min_{t\in N(s)}d_{k-1}(t),\quad d_{k}(e)=0.$
+Algorithm 1 Simple iteration (Dynamic programming)
+1:Hyperparameters: $\epsilon>0$ (tolerance), $\alpha\in(0,1]$ (learning rate), NN (maximum iterations)
+2: Initialize $d_{0}(s)$ for all states $s\in\mathcal{S}$ (e.g., by zeros)
+3:for $i=0$ to NN do
+4:  $d_{i+1}(e)\leftarrow 0$
+5:  for each state $s\neq e$ do
+6:   $d_{i+1}(s)\leftarrow\alpha(1+\min\limits_{a}d_{i}(a\circ s))+(1-\alpha)d_{i}(s)$
+7:  end for
+8:  $c_{i}\leftarrow\mathrm{corr}(\boldsymbol{d}_{i},\boldsymbol{d})$
+9:  if $\|\boldsymbol{d}_{i}-\boldsymbol{d}\|<\epsilon$ then
+10:   break
+11:  end if
+12:end for
+13:Output: Number of iterations $i$, vector of correlations $\boldsymbol{c}$
+Examples of solving Bellman equation.
+Initialization dependence.
+Let us illustrate solving Bellman equation on Cayley graphs and demonstrate some unexpected dependence on initializations.
+We observe that if initialization is positive convergence takes not more than diameter steps (it is no longer true for initialization with negative values).
+However convergence might not be monotonic - depending on initialization, it might happen that first solution deviates from the final result, and only after several steps, begins to converge again.
+For small graphs such as Pocket Cube or LRX with $n\leqslant 10$ the true distances $d(s)$ can be easily calculated via BFS for each state.
+Therefore, we can use these true distances as stopping criteria: when $\|d_{n}(s)-d(s)\|<\varepsilon$, we are finished.
+In addition, we measure performance on each iteration by calculating the correlations between $d(s)$ and $d_{k}(s)$.
+The main question of interest here was how the initialization influences the performance.
+That is important to understand in view of our proposal to use initialization coming from the diffusion distance.
+We tried different initializations $d_{0}(s)$:
+  • zero initialization $d_{0}(s)=0$ for all $s$
+  • initialize each of $m$ layers by a specific value (e.g., first $k$ layers by true distances, last mkm-k layers by zeros, $1\leqslant k<m$)
+  • Manhattan distance to the target: $d_{0}(s)=\sum|s_{k}-k|$
+  • Hamming distance to the target: $d_{0}(s)=|\{k\colon s_{k}\neq k\}|$
+  • random integers form 0 to $m$ or random values from $\mathcal{N}(0,\sigma^{2})$
+  • outputs of a model predicting diffusion distances (see next subsection or formula (10))
+Figure 3. Pearson correlation between $d_{i}(s)$ and $d(s)$ on each iteration of the DP algorithm, $\epsilon=10^{-3}$ (Pocket Cube)
+These experiments revealed interesting phenomena presented in figure 3.
+For some specific initialization the performance first degrades, and only after that achieves the convergence to the true distance which is guaranteed by the general theory.
+The number of steps for convergence does not exceed the diameter of the graph if initialization is positive, however, it can be larger in presence of negative values.
+Initialization coming from the machine learning model (”catboost”) trained on diffusion distance (same as ”warm-up step” in our proposal) improves the convergence, though not in a radical manner.
+Deep Q-learning algorithm.
+Tabular methods like DP or Q-learning are not applicable in situations where the graph is too large.
+A family of deep Q-learning methods like [Mnih13] emerged to resolve this issue.
+Typically, these algorithms leverage a neural network to predict targets like right-hand sides of 4 or 2:
+$f_{\boldsymbol{\theta}}(s)=r+\max_{a}Q_{\boldsymbol{\theta}}(s,a).$
+The algorithm 2 is a deep Q-learning method designed to solve the Bellman equation 2 (here we assume that $s$ is an embedding of a graph node suitable for input of the neural network $f_{\boldsymbol{\theta}}(s)$).
+Typically, the choice of the training subset of nodes is made by walking over the graph, combining both random steps and steps guided by the model.
+After the training is completed, the output of the model $f_{\boldsymbol{\theta}}(s)$ should hopefully serve as a good approximation to the true distance from $s$ to $e$.
+Algorithm 2 Deep Q-learning
+1:Hyperparameters: NN (number of epochs), $M$ (subset size)
+2: Initialize weights $\boldsymbol{\theta}$ of the neural network $f_{\boldsymbol{\theta}}(s)$
+3:for $i=0$ to NN do
+4:  select a random subset of nodes $S_{i}$, $|S_{i}|=M$, $e\notin S_{i}$
+5:  for each state $s\in S_{i}$ do
+6:   $d_{i}(s)\leftarrow 1+\min\limits_{t\in N(s)}f_{\boldsymbol{\theta}}(t)$
+7:  end for
+8:  calculate MSE loss $\mathcal{L}(\boldsymbol{\theta})=\frac{1}{M}\sum\limits_{s\in S_{i}}\big(d_{i}(s)-f_{\boldsymbol{\theta}}(s)\big)^{2}$
+9:  calculate $\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$ and do the gradient step
+10:end for
 Graph path finding.
-To find a path from a given node $g$ to the selected node $e$, one needs to run a greedy search algorithm.
-Starting from node $g$, the algorithm inspects all its neighbors, evaluates the distance function for each of them, and then chooses the node with the minimum value of the distance function, moving to that node.
-The process is repeated until the destination node is found or the step limit is exceeded.
-If the distance function accurately reflects the true distance on the graph (i.e., the length of the shortest path), this algorithm will find the shortest path in an optimal way, performing the least number of operations.
-However, in the case of deep Q-learning, the predictions of the neural network are typically not very precise, and the greedy algorithm can get stuck in local minima.
-To overcome this challenge, more advanced graph search algorithms, such as beam search or A-star, are used.
-This will be discussed further below.
-2.2. Diffusion distance pathfinding - CayleyPy-1
-Let us briefly remind the novel pathfinding approach proposed in our previous paper [Chervov25].
-Its efficiency has been demonstrated on the Rubik’s cube where it significantly outperforms all possible competitors providing shorter solutions than any existing approach for 3x3x3,4x4x4,5x5x5 cubes.
-It is also more computationally efficient and in some sense it is more simple.
-The idea of the approach is to work not with the true distance on the graph, but with the diffusion distance which might be thought as a length of the random path.
+To find a path from a given node $s$ to the target $e$, one needs to run a greedy search algorithm, which, starting from node $s$, inspects all its neighbors and evaluates distance function for them, then chooses the node with the minimum value of the distance function and makes a move to that node.
+The process is repeated until the number of the destination node found at the limit of the step is exceeded.
+If the distance function is the true distance on the graph (the shortest path length), such an algorithm finds the shortest path optimally (least possible operations).
+However, the predictions of the neural network $f_{\boldsymbol{\theta}}(s)$ from the deep Q-learning algorithm 2, are typically not very precise, and the greedy algorithm will be stuck in local minima.
+To overcome that difficulty, one used more advanced graph search algorithms like beam search, A-star, etc.
+A guided beam search approach is described in algorithm 3; on each iteration it maintains $W$ states with lowest estimations to the target state.
+If algorithm 3 can find the path to the target in NN iterations, it returns the length of this path.
+2.2. Diffusion distance pathfinding (CayleyPy-1)
+Let us briefly remind the alternative pathfinding approach proposed in our previous paper [Chervov25].
+Its efficiency has been demonstrated on the Rubik’s cube, where significantly outperforms all possible competitors, providing shorter solutions than any existing approach for $3\times 3\times 3$, $4\times 4\times 4$, $5\times 5\times 5$ cubes.
+In some sense, it is simpler and more computationally efficient.
+Algorithm 3 Guided beam search pathfinding
+1:Hyperparameters: $s_{0}$ (starting state), $M$ (step limit), $f(s)$ (distance estimator), $N(s)$ (neighbors of $s$), $W$ (beam width)
+[…]
+3:for $i=0$ to $M$ do
+4:  $B^{\prime}\leftarrow\varnothing$
+5:  for each state $s\in B$ do
+6:   if $s=e$ then
+7:    return $i$
+8:   end if
+9:   $B^{\prime}\leftarrow B^{\prime}\cup N(s)$
+10:  end for
+11:  $B^{\prime}\leftarrow B^{\prime}\setminus B$
+12:  sort states $s\in B^{\prime}$ by the distance estimations:
+$f(s_{(1)})\leq f(s_{(2)})\leq\ldots\leq f(s_{(|B^{\prime}|)})$
+13:  $B\leftarrow\{s_{(1)},\ldots,s_{(\min\{W,|B^{\prime}|\})}\}$
+14:end for
+The idea is to work not with the true distance on the graph but with the diffusion distance, which is the length of the random path.
+The rigorous definition of the diffusion distance is given by the formula (10) in the supplementary section.
 The diffusion distance is easy to estimate by the number of steps of random walks.
-Generation of random walks is much more fast than computations of the neural networks predictions (the only way to generate target in Q-learning), thus generation of the training data is much more efficient than in Q-learning approach, though the training data might be of weaker quality.
-The assumption is that for each node of a graphs there is feature vector attached.
-For example for permutation groups it is a vector representing a permutation $(p_{0},p_{1},p_{2},...p_{n-1}),p:i\to p_{i}$ We also assume that there is a selected node (e.g.
-identity of the group) where we need to find a path from any other node.
-Core of the approach consists of the three steps:
+Generating random walks is much faster than computing neural network predictions.
+Thus, the training data generation is much more efficient than in Q-learning approaches, though the training data might be of poorer quality.
+The core of our approach consists of three steps:
   1. (1)
-    Generation of training set by random walks.
-Generate $N$ random walks trajectories all starting from a selected node $e$ (identity of a group).
+    Generation of the training set via random walks.
+Generate $N$ random walk trajectories starting from the selected node $e$.
 Each random walk trajectory consists of up to $K_{\text{max}}$ steps, where $N$ and $K_{\text{max}}$ are integer parameters of the method.
-For some nodes encountered during the random walks, we store a set of pairs $(v,k)$, where $v$ represents the vector corresponding to the node and $k$ is the number of steps required to reach it via the random walk.
+If a random walk visits node $s$ for the first time at step $k$, we store the pair $(s,k)$ in our training set.
   2. (2)
-    Generation of training set by random walks.
-The generated set of pairs $(v,k)$ serves as the training set for the neural network.
-Specifically, $v$ serves as the ’feature vector’ (the input for the neural network), and $k$ represents the ’target’ (the output that the network needs to predict).
-Thus, the neural network’s predictions for a given node $v$ estimate the diffusion distance from $v$ to the selected destination node.
+    Training a model.
+The generated set of pairs $(s,k)$ serves as the training data for a supervised machine learning model $f(s)$, typically a neural network.
+For a given node $s$ this model predicts the estimated diffusion distance from $s$ to the target $e$.
   3. (3)
-    Graph search guided by neural network.
-Beam search.
-This step finds a path from a given node to the destination node.
-The neural network provides heuristics on where to make the next steps, while the graph pathfinding technique compensates for any possible incorrectness in the neural network predictions.
-The beam search pathfinding method is quite simple, but has proven to be the most effective for us and works as follows.
-Fix a positive integer $W$ a parameter known as the ”beam width” (or ”beam size”).
-Starting from a given node, we take all its neighboring nodes and compute the neural network predictions for all of them.
-We then select the $W$ nodes that are closest to the destination according to the neural network (i.e., the predictions have smaller values).
-For these selected $W$ nodes, we take their neighbors, drop duplicates, and again compute the neural network predictions, choosing the top $W$ nodes with the best (i.e., minimal) predictions.
-This process is repeated until the destination node is found (or limit of steps is exceeded).
-The last step is common for all methods - based on Q-learning or diffusion distance.
-One uses a neural network predictions as heuristic function and uses graph search algorithms to find a path.
-We are strongly in favor of beam search, while DeepCube team [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] uses A-star algorithm and develops its modifications.
-2.3. Beam search effectivity - toy model of research community
-To some extent, the success of the proposed scheme is due to beam search and our efficient implementation supporting large beams.
+    Guided beam search.
+Here we apply the algorithm 3 to find a path from a given state $s$ to the target node $e$.
+The model $f(s)$ trained on the previous step provides heuristics on where to make the next steps, while the beam search technique compensates for any possible incorrectness in the predictions of the neural network.
+Having fixed a positive integer $W$ (”beam width” or ”beam size”), on each step we select $W$ nodes that are closest to the destination according to the neural network.
+Then we check $f(s)$ for all neighbors of these best $W$ states, drop duplicates, and repeat the process until the destination node is found or the step limit is exceeded.
+The last step is common for any method, whether it is based on Q-learning or diffusion distance.
+One uses a neural network predictions as a heuristic function and graph search algorithms to find a path.
+We are strongly in favor of beam search since it is quite simple, but has proven to be the most effective for us, while the DeepCube team [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] uses the A-star algorithm and develops its modifications.
+A simple but useful modification of the diffusion distance method leverages non-backtracking which is applied during steps (1) and (3).
+On both occasions, we store the history states visited in the $m$ previous steps to avoid returning to them.
+The history of non-backtracking random walks contains $O(mN)$ states, and all trajectories are not allowed to visit any of these states and their neighbors in the next step.
+For non-backtracking beam search we do not include the states from the history on line 9 of the Algorithm 3.
+2.3. Beam search effectivity. Toy model of research community
+To some extent, the success of the proposed above is due to beam search and our efficient implementation supporting large beams.
 We started with a modified greedy search, then moved to Metropolis and A-star, before finally finding relief in beam search.
@@
 Greedy search can get stuck in them, whereas beam search can bypass them entirely if the beam size is larger than the depth of the local minima.
-It’s similar to how an elephant wouldn’t even notice small dips in the terrain, while for an ant, they could be insurmountable obstacles.
-Another important aspect of beam search is that increasing the beam size not only linearly increases computation but also linearly increases memory consumption.
-On one hand, typically memory - not computational time - is the main bottleneck of the method.
-On the other hand, having the largest possible beam size is crucial for its performance.
+It is similar to how an elephant wouldn’t even notice small dips in the terrain, while for an ants, they could be an insurmountable obstacles.
+Another important aspect of beam search is that increasing the beam size increases both computation and memory consumption linearly.
+Typically, memory – not computation time – is the main bottleneck.
+However, maximizing the beam size is crucial for performance.
 Moreover, there is no simple way to reproduce the results of beam search with a beam size of $B+1$ using any algorithm that only consumes memory equivalent to beam size $B$.
-To explain this point, we use an analogy with the work of a research community.
+To explain this point, we will use the analogy with the work of the research community.
 Imagine two researchers working together on a creative problem (a beam size of 2).
-The key factor is the exchange of ideas—if one researcher makes a breakthrough, they share it with the other, allowing both to continue their search from the breakthrough point.
+The key point is the exchange of ideas: if one researcher makes a breakthrough, they share it with the other, allowing both to continue their search from the breakthrough point.
 Without this exchange, the second researcher might get stuck in a dead end.
-This sharing of ideas helps avoid wasting time on unproductive directions.
-The mechanism described above explains why increasing the beam width is so effective—it acts as an exchange of ideas, focusing on breakthroughs rather than following unproductive paths.
+This sharing of ideas helps avoid wasting time in unproductive directions.
+The mechanism described above explains why increasing the beam width is so effective: it acts as an exchange of ideas, focusing on breakthroughs rather than following unproductive paths.
 Essentially, this is what beam search does: it selects the top nodes across the entire beam neighborhood rather than picking the best few for each individual node.
@@
 A kind of natural selection process.
-From the above, it is clear that the size of a research community matters.
-Reducing the community size by half will not simply cut the research output in half — it will decrease it exponentially.
-That aspect is different from the more routine work - cutting it twice output will be cut twice.
+These observations show that community size has a disproportionately large impact on research productivity.
+Halving the number of researchers does not simply halve the output — it causes a super-linear (often exponential) decline.
+In contrast, for routine tasks, a 50% reduction in resources usually yields about a 50% reduction in output.
 The same principle applies to beam search: the beam size is crucial, and there is no simple way to achieve the performance of a beam size $B$ using a smaller beam.
@@
 Exploring various methods on Cayley graph pathfinding tasks is an interesting direction.
-2.4. Novel method - combining diffusion distance with deep Q-learning
-Rationale.
-Why we need that?
-The diffusion distance approach described above is quite simple and efficient however there is practical and theoretical disadvantage.
-The practical side consists of somewhat surprising phenomena that enlarging the training set (generating more random walks) does not always improve the quality of the model - such stagnation was observed for Rubik’s cube cases in our previous paper and it is even more prominent for LRX graphs.
-It is actually related to the quite clear theoretical fact: diffusion distance is not monotonically dependent on the true distance - one node can be further than another for the true distance, while oppositely ranged for the diffusion distance.
-(That problem is not only for random walks estimation, but for the diffusion distance itself (i.e.
-limit of infinite data - computed theoretically (analysis) ).
-So diffusion distance is easy to estimate, but it is not precise.
-On the other hand DQN and Bellman equation allows to approximate the true distance, but for the price of more heavy computations.
-The idea of the proposal below is to combine less precise but light weight diffusion distance with more heavy and more precise Bellman equation approach.
-And to avoid the sparse reward problem.
+2.4. Combining diffusion distance with deep Q-learning
+The diffusion distance approach described above is quite simple and efficient; however, there are practical and theoretical disadvantages.
+The practical side consists of somewhat surprising phenomena that enlarging the training set (generating more random walks) does not always improve the quality of the model; such stagnation was observed for Rubik’s cube cases in our previous paper, and it is even more prominent for LRX graphs.
+It is actually related to the quite clear theoretical fact: diffusion distance is not monotonic function of the true distance.
+This problem is not only for random walks estimation, but also for the diffusion distance itself (e.g., see computed theoretically analysis).
+So, diffusion distance is easy to estimate, but it is not precise.
+On the other hand, DQN and Bellman’s equation allow us to approximate the true distance, but for the price of more heavy computations.
+The idea of the proposal below merges lightweight diffusion distance with a heavier and more precise Bellman equation approach, and at avoiding the sparse reward problem.
 Our proposal looks as follows.
@@
 Warm-up diffusion distance training.
-Use the diffusion distance approach for preliminary training the neural network.
-I.e.
-apply the steps 1,2 above: generate random walks to create training data, and train neural network to predict number of steps of random work.
-Part 2.
-Modified DQN training.
-The second phase will relies on the Bellman equation strategy with mild modifications, but on somewhat different space exploration strategy comparing to conventional DQN - borrowed from the diffusion distance approach:
-Step 1.
-Generate $N$ non-backtracking random walks trajectories all starting from a selected node $e$ (identity of a group).
-Each random walk trajectory consists of up to $K_{\text{max}}$ steps, where $N$ and $K_{\text{max}}$ are integer parameters of the method.
-Step 2. For the obtained nodes compute the Bellman equation predictions $t(g)=1+\min_{n:~{}neigbours~{}of~{}g}d(n),t(e)=0$ in the standard way.
-Step 3. Clip the predictions by the number of steps of the random work since true distance is always smaller than number of steps. (Clip also predictions negative values).
-Step 4. Run gradient descent for the neural network to minimize the loss function between clipped $t(g)$ (target) and neural network predictions.
-Part 3.
-Graph search algorithm - beam search with neural network as an heuristics.
-After the neural network is trained we utilize the same approach to find a path by beam search and neural network as in our previous paper.
-The advantages of the proposed method are the following: there is no sparse reward problem - since we always start from the selected state, it always allows us to clip overestimated predictions, the warm-up phase allows to avoid long initialization process until when targets obtained by Bellman equation random since network starts from random weights - in our approach targets are always meaningful - first they are steps of random walks, at the second phase neural by itself produces meaningful targets since it was preliminary trained.
-We also suggest to use non backtracking random walks by rather clear reason - number of steps for non-backtracking is more similar to true distance.
-Conceptually that is related to the well-known phenomena - ”Non-backtracking random walks mix faster” [Alon07].
-Benchmarks and analysis.
-Below we present the analysis of that proposal.
-We observe stable improvement of the method above comparing to diffusion distance approach, unfortunately the effect is not so big.
-The first experiment is concerned not large groups, where one can compute by brute force the true graph distance and compare it with the predictions of the neural network.
-We see that for small sizes of the group methods learn practically ideally the distance on a graph, however for large groups correlation becomes not so perfect.
-Additional training by DQN always improves the preliminary warm-up training by diffusion distance procedure.
-$n$ epochs: 5+0 epochs: 50+0 epochs: 5+50 epochs: 50+50
-6 0.955 0.99 0.99 0.99
-8 0.954 0.969 0.98 0.982
-10 0.883 0.923 0.939 0.938
-Table 1.
-n - permutation length.
-Spearman correlation of the true distance (obtained by brute force on million states) vs predictions of the neural networks trained by different methods.
-Conclusion Additional training with DQN improves results.
-Epochs A+B denotes: A - training epochs in warm-up diffusion distance part, B - epochs in DQN part.
-The second experiments concerns larger group $S_{28}$ with LRX generators.
-We train the model and search for path by the beam search.
-The length of the path is the measure of performance.
-Experiment confirms that additional training by DQN improves the results.
-The combinatation of the 30+200 epochs training with 30-warmup and 200 - second stage (DQN) - is one the best which we observed for that case.
-Average length is 402, best is 382 which is not far from the expected ideal length 378.
-(In all experiments we searched path from the same element - conjecturally the longest element for LRX generators).
-For each experiment models for retrained from scratch, thus due to random nature of the random walks models are significantly different.
-Experiment epochs: epochs: epochs: epochs:
-30+200 30+0 100+0 0+230
-1 490 None 446 None
-2 396 658 516 None
-3 382 None None None
-4 408 650 436 None
-5 428 706 1046 None
-6 396 None None 508
-7 426 890 614 None
-8 None 598 None None
-9 392 464 524 None
-10 None 626 452 434
-Median 402 650 516 455.5
-Table 2.
-Lengths of solutions obtained by different approaches.
-$S_{28}$ LRX generators.
-10 experiments to decompose the ”longest” element with expected ideal length 378.
-Conclusion: additional training by DQN improves the results - both in length and probability to find a path.
-Pure DQN training is not performing well.
-The analysis can be found in notebook.
-3. Benchmarks
-Goal - $S_{110}$.
-DeepMind’s success with Go game with $10^{170}$ states - creates a challenge to achieve similar results in the other fields in particular for Cayley graphs, at the moment we are somewhat far from that, but there are plenty ideas not yet incorporated in our CayleyPy project and optimistically such a goal is achievable in near horizons.
-Moreover in the present study we achieve pathfinding for such scale groups, but unfortunately not by fully zero prior knowledge.
-Our zero prior knowledge approaches achieves pathfiding for LRX generators of $S_{n}$ a bit above $30$, but at the moment $40$ is a not achieved despite bunch of neural networks we tried.
-What is surprising that almost all approaches we tried stuck slightly above $n=30$ all in the same region, with small variations.
-It is significantly better than results of computer algebra system GAP which achieves only about $n=20$, which much more lengths and timings.
-3.1. One line of code to get orders of magnitude improvement from: $S_{30}$ to $S_{100}$.
-if action == ’X’ and current-state[0] < current-state[1]: continue The condition used in the beam search part.
-Its logic is rather clear - if the contents of positions 0 and 1 are already sorted - it is logical not to swap them i.e.
-not to apply $X$ generator.
-Surprisingly such a simple conditions is a game changer.
-Inserting it and using it with exactly the same pipelines which performed for around $n=30+$ we are able to find paths for $n=100$ and (apparently more - just stopped by 16G RAM limit).
-So the size of graphs changes from $10^{3}0+$ to $10^{1}50+$.
-It works the same for all models: perceptrons, gradient boostings, convolutional neural networks.
-The challenge is how to avoid such prior knowledge method and extend it to other generators.
-It should also be noted that this condition prevents to find optimal paths. And for small sizes of $n\leq 15$ it typically leads to a bit longer paths, but for higher values of $n$ it is a relief.
-Examples of lengths found by that method are presented at table 3 below:
-$S_{n}$ 100 90 80 70 60
-length 16632 33459 7510 8115 3650
-ideal length 4950 4005 3160 2415 1770
-Time 8h 43m 3h 57m 1h 5m 52m 3h
-Nb version 255 242 241 233 -
-Table 3. n - permutation length. Lengths of solutions for different $n$ obtained by ML model combined followed by beam search with ”X-condition”. (Not completely zero prior knowledge approach).
-Computation for $n=100$ took 8h 43m 53s GPU P100.
-Beam width ${2}^{20}$, 30 epochs of training.
-( Notebook version 255), for other cases we used beam size not more than ${2}^{18}$, that is why for $n=90$ length is much bigger.
-The data for $n=60$ is obtained by the gradient boosting model, with slightly better lengths, but more timing, in the other cases single layer MLP model used.
-In all cases we decompose the conjecturally longest element of expected ideal length $n(n-1)/2$.
-3.2. ”Overcome the GAP” with zero prior knowledge AI.
-We use LRX Cayley graph of $S_{n}$ and its conjecturally longest element of expected length $n(n-1)/2$ as a benchmark.
-Below we demonstrate that AI methods significantly outperform classical methods of computer algebra system GAP.
-$n$ GAP Ideal AI
-Length Length n(n-1)/2 Length
+Apply diffusion distance approach for pre-training the neural network, i.e.
+steps (1) and (2) above: generate training data via random walks, and train a neural network $f_{\boldsymbol{\theta}}(s)$ to predict the number of steps of the random walk.
+Part 2. Modified DQN training. We use the deep Q-learning algorithm 2 with two enhancements:
+  • the random subset of nodes consist of all states visited by $N$ random walks of size $K_{\max}$;
+  • the result of the Bellman update $1+\min\limits_{t\in N(s)}f_{\boldsymbol{\theta}}(t)$ is clipped to the interval $[0,k]$ where $k$ is the time step when the state $s$ was visited for the first time.
+The pseudo code for this modified version of DQN is presented in Algorithm 4.
+Part 3. Guided beam search. Here we apply step (3) from the diffusion distance approach described above (Algorithm 3).
+Algorithm 4 Modified deep Q-learning
+1:Hyperparameters: $M$ (number of epochs), NN (number of random walks), $K_{\max}$ (length of random walks), $f_{\boldsymbol{\theta}}(s)$ (distance estimator)
+2:for $i=0$ to $M$ do
+3:  Generate $N$ random walks of size $K_{\max}$ from the target node $e$
+4:  Collect dataset $\mathcal{D}=\{(s,k)\}$ of states $s$ visited by random walks at time $k$
+5:  for each $(s,k)\in\mathcal{D}$ do
+6:   $d(s)\leftarrow 1+\min\limits_{t\in N(s)}f_{\boldsymbol{\theta}}(t)$ (update based on Bellman equation (2))
+7:   $d(s)\leftarrow\min\{k,\max\{0,d(s)\}\}$ (clipping)
+8:  end for
+9:  Calculate MSE loss $\mathcal{L}(\boldsymbol{\theta})=\frac{1}{|\mathcal{D}|}\sum\limits_{\mathcal{D}}\big(d(s)-f_{\boldsymbol{\theta}}(s)\big)^{2}$
+10:  Calculate $\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$ and do the gradient step
+11:end for
+There are several advantages of the proposed method.
+First, it avoids the sparse reward problem: since we always start from the selected state, we can reliably clip overestimated predictions.
+Second, the warm-up phase helps avoid the long initialization typical for standard DQN, where targets from the Bellman equation are noisy because the network starts with random weights.
+In our approach, the targets are meaningful from the beginning: in the first phase, they come from random walk steps; in the second phase, the network itself produces useful targets, as it has already been pre-trained.
+Finally, we use non-backtracking random walks for a clear reason: the number of steps in such walks better reflects the true distance.
+Conceptually, this relates to the well-known phenomenon ”Non-backtracking random walks mix faster” [Alon07].
+3. Analysis and results
+DeepMind’s success in mastering the game of Go, which has approximately $10^{170}$ possible states, poses an inspiring challenge: achieving similar breakthroughs in other domains, such as Cayley graphs.
+At present, we are still a distance from this goal.
+However, many promising ideas have yet to be implemented in our CayleyPy project, and with continued development, reaching such a milestone may be possible in the near future.
+In this study, we demonstrate successful pathfinding in very large LRX Cayley graphs of $S_{n}$.
+We use its conjecturally longest element (1) of expected length n(n1)/2n(n-1)/2 as a benchmark.
+Our zero prior knowledge approach can find the paths from $\ell_{n}$ to $e$ for $n\leqslant 44$.
+This is a significant improvement over the performance of the GAP computer algebra system, which can only handle up to around $n=20$, and with much longer path lengths and runtimes.
+Introducing some domain knowledge (see Section 3.2 below) drastically improves the performance to $n=100$ or even more.
+3.1. ”Overcome the GAP” with zero prior knowledge AI
+Here we demonstrate that the diffusion distance method significantly outperform classical methods of computer algebra system GAP.
+$n$ GAP length $\frac{n(n-1)}{2}$ DD length
 9 41 36 36
@@
 21 - 210 210
-Table 4. n - permutation length. Comparison of GAP vs AI methods (zero prior knowledge). On the conjecturally longest element of LRX Cayley graph. (Its expected ”ideal” length is $n(n-1)/2$.)
-For $n=20$ GAP timing is 41min 18s, while AI methods can find results much faster.
-For example, simple neural network: notebook version 423: 3m 48s GPU P100.
-GAP benchmarks has been performed in the notebook.
-$n$ Ideal AI
-Length n(n-1)/2 Length
-22 231 231
-23 253 253
-24 276 276
-25 300 300
-26 325 325
-27 351 351
-28 378 378
-29 406 408
-30 435 437
-31 465 478
-32 496 728
-33 528 642
-Table 5. n - permutation length. Comparison of GAP vs AI methods.
-The AI results above are aggregated results from experiment with different models and approaches.
-In particular based on gradient boostings and neural networks, all of them quite stably and perfectly solve cases with $n\leq 30$, but going beyond that requires efforts for all our approaches.It is surprising that all types of models have the problem at the same range of $n$ and that length is not becoming gradually longer and longer, but it stays almost perfect, and then quite sharply pipeline loses ability to find a path at all.
-For some cases we use several reruns - if the path not found via single run, it sometimes be found by multiple relaunches of the pipeline.
-For example $n=32$ has been solved with our modified DQN strategy with 30 warmup epochs and 1000 DQN epochs (notebook), time: 2h 53m 53s GPU P100.
+Table 2. Comparison of GAP and diffusion distance ($D$) method for the conjecturally longest element of LRX Cayley graph with expected length n(n1)/2n(n-1)/2.
+For $n=20$, the GAP timing is 41min 18s, while our methods can find results much faster.
+For example, the diffusion distance method with simple neural network requires only 3m 48s of calculations on GPU P100 (notebook version 423).
+GAP benchmarks have been performed here.
+$n$ $\frac{n(n-1)}{2}$ DD length $n$ $\frac{n(n-1)}{2}$ DD length
+22 231 231 31 465 478
+23 253 253 32 496 508
+24 276 276 33 528 642
+25 300 300 34 561 667
+26 325 325 36 630 834
+27 351 351 38 703 759
+28 378 378 40 780 926
+29 406 408 42 861 909
+30 435 437 44 946 1060
+Table 3. Lengths of solutions produced by diffusion distance ($D$) method for the conjecturally longest element of LRX Cayley graph with expected length n(n1)/2n(n-1)/2.
+The DD length columns in Tables 2 and 3 are aggregated results from experiments with different guiding models based on gradient boostings and neural networks.
+All of them solve cases with $n\leq 28$ quite stably; moreover, after some hyperparameter tuning, they find the shortest paths.
+Going beyond the limit $n=28$ requires some effort.
+For $n\leq 44$ a path can be occasionally found by multiple relaunches of the pipeline, however, its length is no longer optimal.
+3.2. X-trick.
+Surprisingly, adding one line of code allows us to drastically increase the size of solvable graph, from $S_{40}$ to $S_{100}$.
+Namely, add the following condition in line 9 of the beam search algorithm 3: if $s_{0}<s_{1}$ then do not include $X\circ s$ in $N(s)$.
+The logic is straightforward: if the first two elements of the permutation are already sorted, it makes little sense to swap them by applying the generator X=(0 1)X=(0\ 1).
+This X-trick turns out to be a game changer.
+By adding it to the pipelines that succeeded for $n$ about $40$, we were able to find paths for $n=100+$ (and potentially even larger values, as we were limited by the 16GB RAM limit).
+This represents a shift from working with Cayley graphs of size $10^{30+}$ to $10^{150+}$.
+This trick worked universally across model types, including perceptrons, gradient boosting machines, and convolutional neural networks.
+The key challenge now is how to eliminate such prior knowledge and generalize the method to other types of generators.
+However, X-trick may prevent the discovery of optimal paths, especially for small values of $n\leq 15$.
+Examples of lengths found by the diffusion distance method with X-trick are presented in Table 4 below:
+$n$ 100 90 80 70 60
+Solution length 16632 33459 7510 8115 3650
+Diameter (proposed) 4950 4005 3160 2415 1770
+Time 8h 43m 3h 57m 1h 5m 52m 3h
+Notebook version 255 242 241 233 -
+Table 4. Lengths of solutions for different $n$ obtained by ML model followed by beam search with X-trick. (Not completely zero prior knowledge approach).
+Computation for $n=100$ took 8h 43m 53s GPU P100.
+Beam width ${2}^{20}$, 30 epochs of training ( Notebook version 255).
+For other cases we used beam size $\leq{2}^{18}$, that is why for $n=90$ length is much larger.
+The data for $n=60$ is obtained by the gradient boosting model, with slightly better lengths, but more timing, in the other cases single layer MLP model used.
+In all cases we decompose the conjecturally longest element of expected length n(n1)/2n(n-1)/2.
+3.3. Diffusion distance + modified DQN
+We observe that replacing the training stage with the DQN algorithm 4 gives a consistent improvement over the diffusion distance approach; however, the effect is not very large.
+Below we provide two illustrative experiments.
+$n$ epochs: $5+0$ epochs: $50+0$ epochs: $5+50$ epochs: $50+50$
+6 0.955 0.99 0.99 0.99
+8 0.954 0.969 0.98 0.982
+10 0.883 0.923 0.939 0.938
+Table 5.
+Spearman correlation between the true distances and the predictions of the neural network trained for AA epochs during the warm-up diffusion distance phase and $B$ epochs during the DQN phase (epochs: $A+B$).
+Additional training with DQN leads to improved results.
+The first experiment (see Table 5) focuses on relatively small groups, where the true graph distance can be computed by brute force and compared with the neural network’s predictions.
+For small group sizes, the method learns the graph distances almost perfectly.
+However, for larger groups, the correlation between predictions and true distances becomes less accurate.
+Still, additional training using the modified DQN always improves upon the initial warm-up phase based on the diffusion distance procedure.
+Warm-up epochs DQN epochs success rate shortest path median path
+30 200 80% 382 402
+30 0 70% 464 650
+100 0 70% 436 516
+0 230 20% 434 471
+Table 6.
+Path lengths for decomposing the longest element in the LRX Cayley graph of $S_{28}$.
+Results in each row are based on 10 independent runs.
+Additional training with DQN improves both the average path length and the success rate of finding a path.
+Pure DQN training alone performs poorly.
+The second experiment concerns the larger group $S_{28}$ with LRX generators.
+We train the model and use beam search to find a path, where the path length serves as the performance metric.
+The results provided in Table 8 confirm that additional training with DQN improves performance.
+The combination of $30$ warm-up and $200$ DQN epochs produces one of the best outcomes observed in this setting.
+The median (shortest) path length $402$ ($382$) is very close to the expected ideal length of $378$.
+For each run, models were retrained from scratch; thus, due to the randomness of the walks, the resulting models differ significantly.
+The analysis can be found in the Kaggle notebook.
+3.4. Non-backtracking
+We exploit non-backtracking idea in two ways - first we can use non-backtracking random walks, and the second we can use non-backtracking in beam search. The second way shows quite a good effect.
+history size, $m$ success rate median length min length
+2 $40.41\%$ $476$ $380$
+4 $40.10\%$ $474$ $382$
+8 $40.32\%$ $471$ $386$
+16 $37.40\%$ $470$ $382$
+32 $37.47\%$ $470$ $378$
+40 $36.16\%$ $480$ $390$
+48 $37.39\%$ $465$ $380$
+128 $36.45\%$ $476$ $382$
+254 $37.85\%$ $468$ $382$
+400 $37.40\%$ $474$ $384$
+512 $36.58\%$ $476$ $384$
+Table 7. Performance of the $D$ method with non-backtracking random walks ($n=28$).
+Beam Solution Length
+Width Non-backtracking RW
+$2^{10}$ None
+$2^{12}$ 252
+$2^{14}$ 202
+$2^{16}$ 194
+$2^{18}$ 194
+$2^{20}$ 190 (Ideal)
+Table 8. Influence of beam width and non-backtracking random walks for solving the longest element of the LRX Cayley graph $S_{20}$.
+The theoretical reason why non-backtracking is helpful is clear for the random walks.
+Imagine our graph is a tree, then for a non-backtracking random walk, the number of steps coincides with the true distance (length of the shortest path).
+Thus diffusion distance coincides with the standard distance.
+For complicated graphs, this is not the case, but still, it forces the diffusion distance to be closer to true distance and to have less variability.
+Here by variability we mean the situation when the same node can be achieved in a different number of steps.
+This is equivalent to having noise in the training data which makes models less precise.
+Similar for beam search non-backtracking mitigates stagnation in local minimums.
+history size $m$ success rate median length min length
+2 $17.65\%$ $430$ $384$
+4 $17.08\%$ $432$ $390$
+8 $21.37\%$ $440$ $382$
+16 $30.07\%$ $446$ $388$
+32 $38.55\%$ $473$ $384$
+64 $43.29\%$ $495$ $380$
+128 $56.95\%$ $517$ $384$
+256 $60.61\%$ $588$ $384$
+512 $67.33\%$ $582$ $384$
+1024 $9.71\% 66$3663 $382$
+Table 9. Performance of the $D$ method with non-backtracking beam search ($n=28$). The path is found more often as the history size grows.
+Injecting non-backtracking into the beam search stage (Algorithm 3) allows to increase the success rate: the greater is the number of banned states not included in the beam, the higher is the probability of finding a solution (see Table 9).
+On the other hand, the median solution length tends to increase when the history size is large.
+But that cannot be interpreted as performance degradation, rather it means that pipeline can solve more difficult cases, and hence getting large lengths.
 3.5. Architectures and parameter dependence.
-The tests of different architectures and parameters is discussed below.
-The main conclusion that we observe that rather simple models like MLP and gradient boostings currently perform better, than more advanced: transformers, convolutional neural networks etc.
-One exception is one of our transformer models that can find paths up to $n=100$, however it is not zero prior knowledge - it uses features specific to LRX generators.
-Dependence on various parameters was explored in our previous paper  [Chervov25] for the case of the Rubik’s cube group, the conclusions were that beam size is one the most important parameter the length of solution is almost linearly improving on logarithm of beam size, deeper neural networks are better than wider especially, but the effect of depth is better seen on small for higher cubes, we also observed stagnation of performance with increase of training set.
+We observe that rather simple models like MLP and gradient boostings currently perform better than more advanced ones: transformers, CNNs, etc.
+An exception is one of our transformer models that can find paths up to $n=100$, however, it is not a zero prior knowledge model since it uses features specific to LRX generators.
+The dependence on various parameters was explored and discussed in our previous paper  [Chervov25] for the case of the Rubik’s cube group.
+The beam size proved to be one of the most important parameters: the length of solution is almost linearly improving on logarithm of beam size.
+Another observation is that it is more important for a neural network to be deep than wide, however, the effect of depth is better seen for large groups.
+We also noted the performance stagnation as size of the training set increased.
 For the LRX Cayley graph these phenomenons are less strongly expressed, or even negligible.
 Most experiments with deep neural networks have not lead to significant improvements over simple one layer perceptron with 128 neurons.
-Beam size effect is present, in but less expressed (see table below).
-Nonbacktracking random walks have quite strong effect.
-In particular one can train models in just single epoch within a minute for e.g.
-$S_{20}$.
-We actually use improved version of the nonbacktracking based on the following ideas: 1) we forbid not only 1 step in history but typically 32 steps (it is a parameter) 2) we consider a bunch of trajectories and ban states for all of them at once 3) we ban not only visited states but their neighbors.
-The price for that is slowing down the generation, but that is not so important since gradient descent takes much longer time, and overall there is no significant slow down.
-Beam Solution Length Solution Length
-Width Nonbacktracking RW Simple RW
-2**10 None None
-2**12 252 None
-2**14 202 None
-2**16 194 None
-2**18 194 None
-2**20 190 (Ideal) None
-Table 6. Without non-backtracking system cannot perform well on single epoch training. Effect of beam size is clear, but not so strong.
-We generate 10000 trajectories at once each of the length 190, by the generalized nonbacktracking procedure above and train the neural network with batch size 1024 just for single epoch.
-Entire time for each computation - both training and beam search are 1-5 minutes, experiments can be found in versions 427-430 of the notebook.
-Theoretical reason why non-backtracking is helpful is clear.
-Imagine our graph is a tree, then even for non-backtracking random walk number of steps coincide with the true distance (length of the shortest path).
-For complicated graphs it is not so, but still it more close to true distance and has less variability.
-I.e.
-the situation when the same node can be achieved in different number of steps - the fact which equivalent to having a noise in the training data and which makes models less precise.
+In particular, one can train models in just a single epoch within a minute, e.g., for $S_{20}$.
+The beam size effect is still present but not as pronounced (see table 8).
 CNN.
-We also analyse convolutional neural networks - however they have slower training process.
-They show similar results comparing to perceptrons e.g.
-the length of the solution for $n=16$ is 122 (ideal is 120), for $n=20$ is 216 (ideal is 190).
-The basic version of the architecture uses two 1D CNN layers with Batch Normalization and Dropout to avoid overfitting.
+Convolutional neural networks have demonstrated a slower training process.
+Their results are similar to those of perceptrons, e.g., length of the solution for $n=16$ is 122 (ideal is 120), for $n=20$ is 216 (ideal is 190).
+The basic version of the architecture uses two 1-d CNN layers with batch normalization and dropout to prevent overfitting.
 One of the successful variations is using temporal gated CNN.
@@
 Transformer.
-We also peformed experiments with a transformer model, however it was able to solve only not so big groups like $n\leq 15$.
+We also performed experiments with a transformer model.
+However, it was able to solve only not-so-large groups like $n\leq 15$.
 That model incorporated standard transformer components, including multi-head self-attention, positional encoding, feed-forward layers, and dropout to mitigate overfitting.
@@
 This issue became increasingly severe as the sequence length grew.
-For instance, with size = 9 (corresponding to path length = 107 and ideal = 36) or n-size = 13 (path length = 332, ideal = 78), increasing the projection dimension to 512 or 1024, adding more layers, or training for more epochs failed to address the underlying challenges.
+For instance, with $n=9$ (corresponding to path length = 107 and ideal = 36) or $n=13$ (path length = 332, ideal = 78), increasing the projection dimension to 512 or 1024, adding more layers, or training for more epochs failed to address the underlying challenges.
 Notebooks: 1, 2.
@@
 The features inspired by Hamming distance were used: difference between an element and its position; difference with the left neighbor; difference with the right neighbor.
-A standard multihead attention with one encoder layer and mlp was used to estimate the number of steps required to reach the ordered state.
+A standard multihead attention with one encoder layer and MLP was used to estimate the number of steps required to reach the target state.
 The transformer demonstrated the ability to handle sequences up to $n=15$ without using beam search.
-For larger N, beam search is employed, and the maximum N for which results were obtained was 100.
-The model and features only indirectly depend on N, therefore, when trained on micro batches constructed as random walks from an ordered permutation for N*(N-1)/2 moves, it can be used for different N values, both larger and smaller.
+For larger $n$, beam search is employed, and the maximum $n$ for which results were obtained is 100.
+The model and features only indirectly depend on $n$, therefore, when trained on micro batches constructed as random walks from the identity permutation for n(n1)/2n(n-1)/2 moves, it can be used for different values of $n$, both larger and smaller.
 (Notebook).
-Gradient boostings.
-We also applied CatBoost and XGboost in our diffusion distance pipeline, surprisingly they perform not worse than neural networks (it was not the cases for Rubik’s cube - where boosting were not able to solve 4x4x4, 5x5x5 cubes), and even giving a bit shorter solutions.
+Gradient boosting.
+We also applied CatBoost and XGboost in our diffusion distance pipeline.
+Surprisingly, they perform not worse than neural networks (it was not the cases for Rubik’s cube where boostings were not able to solve $4\times 4\times 4$, $5\times 5\times 5$ cubes), and even produce a bit shorter solutions.
 (Code).
-3.4. Reinforcement learning on ”small graphs”
-At the preliminary stage of the research we performed experiments with RL/dynamical programming methods for small graphs.
-In particular the question was how the initialization influence the performance.
-That is important to understand in view of our proposal to use initialization coming from the diffusion distance.
-These experiments revealed interesting phenomenons presented at figure 2.
-For some specific initialization the performance first degrades, and only after that achieves the convergence to the true distance which is guaranteed by the general theory.
-The number of steps for convergence is not more than diameter of the graph is initialization is positive, however can be larger in presence of negative values.
-Initialization coming from the machine learning model (”catboost”) trained on diffusion distance (same as ”warmup step” in our proposal) improves the convergence, though not in a radical manner.
-Figure 2. Pearson correlation between $d_{i}(s)$ and $d(s)$ on each iteration of the DP algorithm, $\epsilon=10^{-3}$
+More simulation results for various parameters can be found in spreadsheet here, or on google sheets.
 4. Mathematical contributions
-In the section we prove lower and upper bounds for the diameter and for particular elements, describe decomposition of the conjecturally longest element of the desired length $n(n-1)/2$, describe algorithms to decompose elements in LRX generators and formulate several conjectures on the growth, random walks and spectrum of the LRX Cayley graph.
-4.1. LRX Cayley graph
-LRX generators are: L - left cyclic shift, R - right cyclic shift, X - eXchange (flip) of positions 0,1.
-It can be considered as a relative of the TopSpin and Hungarian Ring puzzles (figure 3).
-Despite its simplicicity its structure is not fully understood.
-In particular according to OEIS-A186783 the diameter conjecture $n(n-1)/2$ remains open.
-Figure 1 is a visualization of the LRX Cayley graph for $S_{5}$ which shows its complicated structure.
-More visualizations by the minimal distortion method can be found at notebook, and by originally modified force directed layout here.
-Figure 3. LRX and related mechanical puzzles
+In this section, we prove lower and upper bounds for the diameter and for particular elements, describe the decomposition of the conjecturally longest element of the desired length n(n1)/2n(n-1)/2, describe algorithms to decompose in LRX generators, and formulate several conjectures on the growth, random walks and spectrum of the LRX Cayley graph.
 4.1. Long and the longest element
-Figure 4. Permutations arising from dihedral symmetries - are provebly long elements. The longest is conjecturally one of them. Surprisingly it is not the full inversion.
-We created efficient implementation of the brute force breadth first search algorithm notebook using bit operations over int64 and were able to traverse all the permutations for $n\leq 14$, outperforming OEIS data by one step.
-From that computation we observed that the longest element is unique and have a specific form.
-It can be described as a symmetry of a regular polygon (figure 4), which is orthogonal to 0-1 edge.
-Or explicitly as product of transpositions $(0,1)(n-1,2)(n-2,3)(n-3,3)...$.
-We conjecture that this element is unique longest element and it has length $n(n-1)/2$.
-Surprisingly not the full inversion ($i\to n-i-1$) is the longest.
-As we will prove below actually all the permutations which correspond to symmetries of a regular polygon are also long elements.
+Figure 4. Permutations arising from dihedral symmetries are provably long elements. The longest is conjecturally one of them.
+We created an efficient implementation of the brute force breadth first search algorithm (see this notebook) using bit operations over int64 and were able to traverse all the permutations for $n\leq 15$, outperforming OEIS data by two steps (for $n=15$ computation took several days on Hadoop cluster).
+From that computation we observed that the longest element is unique and has a specific form (1).
+It can be described as a symmetry of a regular polygon (Fig.
+4), which is orthogonal to 0–1 edge.
+We conjecture that this element is the only longest element with length n(n1)/2n(n-1)/2.
+Surprisingly, the full inversion $(n-1,n-2,\ldots,2,1,0)$ is not the longest.
+As we will prove below, actually all the permutations which correspond to symmetries of a regular polygon are also long elements.
 Proposition 4.1.
-The element $(0,1)(n-1,2)(n-2,3)(n-3,3)...$ (conjecturally the longest) has the following explicit decomposition into product of LRX generators:
-$\prod_{i=1}^{[(n-1)/2]}(XL^{(-1)^{(i-1)}})^{i}\prod_{i=([n/2]-1)}^{1}(L^{(-1)^{i+[(n-1)/2]-[n/2]}}X)^{i}R^{[n/2]}$ (4.1)
+Let
+$n>2,\quad m=\Big\lfloor\frac{n}{2}\Big\rfloor,\quad\delta_{n}=[n=2m]=\begin{cases}1,n\text{ \rm{is} even},\\ 0,n\text{ \rm{is} odd}.\end{cases}$
+The element $\ell_{n}=(1\ 0\ n-1\ n-2\ldots 3\ 2)$ has the following explicit decomposition into product of LRX generators:
+$\ell_{n}=\prod_{i=1}^{m-\delta_{n}}\big(XL^{(-1)^{i-1}}\big)^{i}\circ\prod_{i=m-1}^{1}\big(L^{(-1)^{i-\delta_{n}}}X\big)^{i}\circ R^{m}.$ (5)
 The proof will be given elsewhere. The decompositions of that type were first found computationally: notebook.
@@
 The key idea is that cyclic shifts preserve the cyclic order of any given triple of numbers.
-Therefore, for each triple where the order is disrupted, t $X$ is required in the decomposition.
+Therefore, for each triple where the order is disrupted, the move $X$ is required in the decomposition.
 The symmetries of a regular polygon, as described above, correspond precisely to the permutations that maximize changes in cyclic order.
@@
 Proposition 4.2.
-Diameter of LRX graph is larger or equal $n(n-1)/2-n/2-1$
-At first, we reformulate the setup as follows.
+The diameter of the LRX Cayley graph of $S_{n}$ is larger than or equal to n(n1)/2n/21n(n-1)/2-n/2-1.
+First, we reformulate the setup as follows.
 The numbers $1,\ldots,n$ are written in the vertices of a regular $n$-gon $A_{1}A_{2}\ldots A_{n}$, each number being written once.
-At each move, one can either rotate the arrangement by $2\pi/n$ counterclockwise or clockwise, or switch the numbers at $A_{1}$ and $A_{2}$.
+At each move, one can either rotate the arrangement by $2\pi/n$ counterclockwise or clockwise, or swap the numbers at $A_{1}$ and $A_{2}$.
 For two arrangements $\pi_{1},\pi_{2}$ denote by $d(\pi_{1},\pi_{2})$ the minimal number of moves which suffice to get $\pi_{2}$ from $\pi_{1}$.
@@
 Assume that $\pi_{2}$ is obtained from $\pi_{1}$ by several moves.
-Call two numbers $a,b\in\{1,\ldots,n\}$ friends, if they were switched odd number of times.
+Call two numbers $a,b\in\{1,\ldots,n\}$ friends, if they were swapped odd number of times.
 Since for every three numbers $a,b,c$ the orientation of a triangle formed by these numbers changed, there is odd number of pairs of friends between $a,b,c$.
@@
 Thus, the total number of pairs of friends is $n(n-1)/2-|A|\cdot|B|\geqslant n(n-1)/2-n^{2}/4$.
-Thus there were at least as many switches.
-Between every two switches there was a rotation (otherwise we do something useless).
+Hence, there were at least as many swaps.
+Between any two swaps there should be a rotation (otherwise, we are doing something useless).
 Totally, the number of operations is not less than $n(n-1)-n^{2}/2-1=n^{2}/2-n-1$.
@@
 We develop two algorithms which can decompose any permutation in the product of LRX generators.
-For the first we prove complexity bound $n(n-1)/2+3n$, thus bounding diameter from above.
+We prove complexity bound n(n1)/2+3nn(n-1)/2+3n, thus bounding diameter from above.
 The code and tests for the first algorithm can be found in notebook.
@@
 The sign of the change depends on the direction of the elementary transposition.
-Essentially, $x$ represents the current position of the transposition, but if our transposition is $(1,N)$, the change in $x$ is also ±$\pm 1$, even though the position changes from 1 to $N$ or from $N$ to 1.
+Essentially, $x$ represents the current position of the transposition, but if our transposition is $(1,N)$, the change in $x$ is also $\pm 1$, even though the position changes from 1 to $N$ or from $N$ to 1.
 4) Next, we attempt to restore the position of $a_{2}$ and compute the sequence of elementary transpositions that move $a_{2}$ to the $a_{2}$-th position.
@@
 Complexity estimation.
-First, when decomposing a cycle into transpositions, we do not shift anything by more than 1, except for elements that have already been affected by the cycle.
+First, when decomposing a cycle into transpositions, we do not shift anything by more than 1, except for elements already affected by the cycle.
 Second, when moving an element from position $i$ (or $i\pm 1$) to position $\sigma(i)$, we perform a sequence of at most $\text{dist}(i,\sigma(i))+1$ transpositions.
@@
 We also develop a second algorithm: notebook.
-Currently it is our top performing algorithm from the practical tests, we expect its complexity is bounded by $n(n-1)/2+n/2$, however it is not yet proved.
+Currently, it is our top performing algorithm from the practical tests, we expect its complexity is bounded by n(n1)/2+n/2n(n-1)/2+n/2, however, it is not yet proved.
 It might be expected that optimal algorithm with polynomial complexity can be developed, but it is not achieved yet.
@@
 [Diaconis77, Diaconis88, Diaconis17] where for Coxeter generators growth and some other statistics were shown to follow Gaussian normal law in a limit.
-New point of our conjecture - appearance of asymmetric Gumbel distribution.
-Such asymmetry is quite natural - for the randomly chosen generators growth will be highly asymmetric with exponential growth and sharp abrupt (Rubik’s cube growth is a prototypical example), while for Coxeter generators which are nearly commutative - growth is symmetric Gaussian close to commutative case.
-Slight asymmetry of growth for LRX generators indicates its intermediate nature between ”exponential growth”/”commutative-like Gaussian growth” cases.
-The growth - Gumbel. The growth distribution of the LRX graph follows the Gumbel distribution for large $n$.
+A new point of our conjecture is the appearance of asymmetric Gumbel distribution.
+Such asymmetry is quite natural since for the randomly chosen generators growth will be highly asymmetric with exponential growth and sharp abrupt (Rubik’s cube growth is a prototypical example), while for nearly commutative Coxeter generators the growth is symmetric Gaussian.
+A slight asymmetry of growth for LRX generators indicates its intermediate nature between ”exponential growth” and ”commutative-like Gaussian growth” cases.
+The growth distribution of the LRX graph seems to follow the Gumbel distribution for large $n$. The detailed analysis is provided in supplementary section 6.1.
 Figure 5. Spectrum distribution of the LRX graphs.
 The analysis of the spectrum of LRX graphs is performed in notebook. From the figure above it is natural to expect that spectrum tends to uniform distribution.
-We have computed all possible shortest paths from the conjecturally longest element to the identity by methods of the dynamical programming (notebook ).
-And after that analyzed averaged trajectories (notebook) - which are analogs of random sorting networks [Virág].
+We have computed all possible shortest paths from the conjecturally longest element to the identity by methods of the dynamical programming (notebook .
+And after that analyzed averaged trajectories (notebook) which are analogs of random sorting networks [Virág].
 The figure 6 below represents the results.
@@
 Figure 6. Analogs of ”random sorting networks” for LRX. Trajectories of individual elements under average of all possible shortest trajectories between longest node and identity.
-P.Diacons and L.Soloff-Coste studied random walks on LRX Cayley [Diaconis93].
-They estimated that mixing time is between $n^{2}$ and $n^{3}log(n)$.
-We performed numerical experiments to make more precise conjecture (notebooks 1,2) unfortunately the achieved range of $n$ was not enough to draw conclusions, however it seems that mixing time is more close to $n^{3}log(n)$, moreover even for non-backtracking random walks, which is a bit surprising since they typically mix faster.
-Finally we considered the Schreier coset graph for the action of $S_{2n}$ on binary strings with $n$ zeros and $n$ ones.
-(Analogue of $Gr(n,2n)$ over the ”field with one element”).
-Efficient code which stores such strings in int64 allowed us to completely traverse the graph up to $n=42$ and compute its growth.
-The analysis of the data presented in the notebook.
-Based on that it is tempting to conjecture that the diameters of such graphs are between $n^{2}/6$ and $n^{2}/5$, while growth again can be approximated by the Guimbel distribution.
 5. Code availability
@@
 The link above leads to one of the three public Kaggle challenges which we created to stimulate the research and interplay between artificial intelligence and mathematical communities, the other two: LRX discover math and God’s algorithm, LRX-binary: in search of God’s number.
-Kaggle infrastructure provides a convenient way to benchmark and compare different algorithms - just making the submissions to these challenges.
+Kaggle infrastructure provides a convenient way to benchmark and compare different algorithms by just making the submissions to these challenges.
 The code, the data and discussions are stored at the same place.
-Code can be executed for free on Kaggle cloud servers.
-The first challenge asks to decompose the conjecturally longest elements and to find their shortest decompositions - participants already achieved the length $n(n-1)/2$ an so if believe in the conjecture the optimum is found, if we do not believe - not.
-The second challenges asks to decompose random permutations and aims to find optimal algorithms (at the moment of writing optimum is not achieved), the third one asks the same but for the binary strings acted by LRX generators.
+The code can be executed for free on Kaggle cloud servers.
+The first challenge is to decompose the conjecturally longest elements and to find their shortest decompositions.
+The participants have already achieved the length n(n1)/2n(n-1)/2; thus, if we believe the conjecture is true, the optimum is found.
+The second challenge is to decompose random permutations and aims to find optimal algorithms (at the moment of writing optimum is not achieved).
+The third one asks the same but for the binary strings acted by LRX generators.
 Acknowledgements
@@
 is deeply grateful to M.
-Douglas for his interest in this work, engaging discussions, and invitation to present preliminary results at the Harvard CMSA program on ”Mathematics and Machine Learning” in Fall 2024, to M.
+Douglas for his interest in this work, engaging discussions, and invitation to present preliminary results at the Harvard CMSA program on ”Mathematics and Machine Learning” in Fall 2024; to M.
 Gromov, S.
 Nechaev, and V.
-Rubtsov for their invitation to make a talk (video) at ”Representations, Probability, and Beyond: A Journey into Anatoly Vershik World” at IHES, as well as for stimulating discussions.
+Rubtsov for their invitation to give a talk (video) at ”Representations, Probability, and Beyond: A Journey into Anatoly Vershik World” at IHES, as well as for stimulating discussions.
 A.C.
 is grateful to J.
-Mitchel for involving into the Kaggle Santa 2023 challenge, from which this project originated and to M.Kontsevich, Y.Soibelman, S.Gukov, A.
+Mitchel for involving into the Kaggle Santa 2023 challenge, from which this project originated, and to M.Kontsevich, Y.Soibelman, S.Gukov, A.
 Hayat, T.
@@
 Melies, I.
-Vlassopoulos, F.Khafizov, A.Zinovyev, H.Isambert for the discussions, interest and comments, to his wife A.Chervova and daugther K.Chervova for support, understanding and help with computational experiments.
+Vlassopoulos, F.Khafizov, A.Zinovyev, H.Isambert, T.
+Rokicki for the discussions, interest and comments, to his wife A.Chervova and daugther K.Chervova for support, understanding and help with computational experiments.
 We are deeply grateful to the many colleagues who have contributed to the CayleyPy project at various stages of its development, including: N.Bukhal, J.Naghiev, A.Lenin, E.Uryvanov, A.